### v2

# Counterfactual Sensitivity in Quantitative Spatial Models

Abstract
Counterfactuals in quantitative spatial models are functions of the current state of the world and the model parameters. Current practice treats the current state of the world as perfectly observed, but there is good reason to believe that it is measured with error. This paper provides tools for quantifying uncertainty about counterfactuals when the current state of the world is measured with error. I recommend an empirical Bayes approach to uncertainty quantification, which is both practical and theoretically justified. I apply the proposed method to the applications in Adao, Costinot, and Donaldson (2017) and Allen and Arkolakis (2022) and find non-trivial uncertainty about counterfactuals.
1 Introduction
Economists use quantitative spatial models to answer counterfactual questions. For example, what is the effect on welfare when trade costs between a set of countries decrease because of an infrastructure investment or a trade agreement? How will expenditure shares change if some countries obtain a novel technology which increases their production efficiency? These counterfactual questions generally have the same structure: what will happen to the endogenous variables when the exogenous variables change in some specific way?
These quantitative spatial models are assumed to exactly match the data, and allow us to write the counterfactual change in an endogenous variable of interest as a function solely of the observables. However, the observables are often noisily measured, which causes meaningful uncertainty about counterfactual predictions. Furthermore, this generates a novel measurement error problem because the object of interest is a function of the true realized values of the observables rather than a function of the correctly measured distribution of the observables, as traditionally considered in the econometrics literature on measurement error.
In this paper, to provide tools for quantifying uncertainty in this novel measurement error setting, I propose an empirical Bayes (EB) approach. This approach takes as inputs a model for the measurement error and a prior on the true data and allows researchers to incorporate economic knowledge through the prior. I outline how to calibrate the parameters of the prior using the data and find the posterior distribution for the counterfactual change in an endogenous variable of interest. From this posterior we can obtain credible sets. In the appendix I also develop an adversarial alternative to the EB approach, which derives worst-case bounds for the counterfactual under researcher-selected bounds on the measurement error.
To fix ideas, consider as an example the canonical Armington model (Armington, 1969). Under this model, we can write the proportional change in welfare as result of a proportional shock to trade costs as a function of baseline bilateral trade flows, baseline incomes and the trade elasticity. We do not know the trade elasticity and hence estimate it using the observed baseline bilateral trade flows. The question I will aim to answer is how mismeasurement in the bilateral trade flows generates uncertainty around the predicted change in welfare.
As illustrated by the Armington model, the structural parameters of the model are often estimated using the noisily measured data. Hence, in the presence of measurement error we must account for estimation error, the direct effect of mismeasurement, and the indirect effect of mismeasurement through the estimation procedure. To address this, we aim to cover the “oracle” confidence interval that accounts for estimation error, based on the correctly-measured data, with prescribed probability.
The EB approach requires researchers to specify a model for the measurement error and a prior on the true data. I provide a widely applicable default approach, which proposes default choices for the measurement error and the prior on the true data when our object of interest is a function of a set of noisily measured positive flows between locations. Specifically, I recommend a normal-normal model where the prior is centered on a gravity model. This default approach can be applied out-of-the-box to most quantitative spatial models, but can also easily be adapted to specific settings.
To illustrate the impact of accounting for measurement error, I revisit the applications in Adao, Costinot, and Donaldson (2017) and Allen and Arkolakis (2022). For the counterfactual analysis in Adao, Costinot, and Donaldson (2017), which considers welfare effects from China joining the WTO, I model measurement error in the bilateral trade flows. I use the widely applicable default approach and calibrate the parameters of the prior and the measurement error model using the mirror trade dataset of Linsi, Burgoon, and Mügge (2023). This dataset collects the reported bilateral trade flows by both the exporter and importer, which I interpret as two noisy measures of the true bilateral trade flow. I plot the change in China’s welfare from 1996 to 2011, and add credible sets that account for measurement error and estimation error. These sets are wide, and for multiple time periods it becomes ambiguous whether joining the WTO had a positive or negative effect on China’s welfare.
For the application from Allen and Arkolakis (2022), the counterfactual question of interest is which highway links in the United States have the highest return on investment, and are hence most worth improving. I model mismeasurement in traffic flows, and again use the widely applicable default approach. I calibrate the measurement error model based on estimates from Musunuru and Porter (2019). I calculate the credible sets that account for measurement error and estimation error for the three links with the highest return on investment. These sets are large, and when estimation error is considered we lose the ability to clearly rank the links.
This paper contributes to the small literature on improving counterfactual calculations in quantitative spatial economics (Ossa, 2015; Adao, Costinot, and Donaldson, 2017; Adão, Costinot, and Donaldson, 2023). The most relevant paper in this literature is Dingel and Tintelnot (2020), which studies calibration procedures in granular settings. It considers models which in addition to assuming that the data are perfectly explainable, also assume there is a continuum of agents. Then, in the case where there are only a handful of observations, individual idiosyncrasies do not wash out and are incorporated in the model, which causes overfitting and poor performance out-of-sample. I focus on the complementary question of uncertainty quantification due to measurement and estimation error, which arises even in non-granular settings.
I make a novel connection between the econometrics literature of measurement error modeling and the literatures on quantitative spatial economics. As is argued in Goes (2023) and Linsi, Burgoon, and Mügge (2023), measurement error is prevalent in the data underlying quantitative spatial models. The most closely related strand of measurement error literature is that on nonseparable error models (Matzkin, 2003; Chesher, 2003; Hoderlein and Mammen, 2007; Matzkin, 2008; Schennach, 2016). The key difference between the current setting and conventional measurement error settings is that the object of interest directly depends on the correctly measured data observations rather than on the correctly measured distribution of the data. I discuss this further in Section 3.1.1 below.
The rest of this paper is organized as follows. The next section introduces the notation and the setting I consider. Section 3 discusses measurement error in quantitative spatial models and introduces my empirical Bayes method for uncertainty quantification. Section 4 applies my method to the application in Adao, Costinot, and Donaldson (2017). Section 5 considers uncertainty quantification in the economic geography model from Allen and Arkolakis (2022). Section 6 concludes.
2 Counterfactuals in Quantitative Spatial Models
This section introduces the notation I will be using and discusses the assumptions that commonly underly counterfactual analyses in quantitative spatial models.
2.1 Notation and Key Assumption
To state the key assumption on quantitative spatial models, I begin by discussing the setting without measurement error. Consider an equilibrium, denoted by $\left(X_{O},X_{U},N_{O},N_{U}\right)$. Here, $X_{O}\in\mathcal{X}_{O}\subseteq\mathbb{R}^{d_{X,O}}$ are exogenous observables, $X_{U}\in\mathcal{X}_{U}\subseteq\mathbb{R}^{d_{X,U}}$ are exogenous unobservables, $N_{O}\in\mathcal{N}_{O}\subseteq\mathbb{R}^{d_{N,O}}$ are endogenous observables, and $N_{U}\in\mathcal{N}_{U}\subseteq\mathbb{R}^{d_{N,U}}$ are endogenous unobservables. Exogenous variables are fundamentals that the model takes as given. Endogenous variables are the variables researchers try to explain, and are determined within the model. [^1] One might argue that “external” and “internal” are less ambiguous terms than “exogenous” and “endogenous”, respectively (LeRoy, 2004). I opt for the latter because of the convention in the literature on quantitative spatial models. These equilibrium variables are connected to each other through the equilibrium conditions. These are, for a given parameter $\theta\in\Theta\subseteq\mathbb{R}^{d_{\theta}}$, given by
$f\left(X_{O},X_{U},N_{O},N_{U};\theta\right)=\boldsymbol{0},$ (1)
for some function $f:\mathcal{X}_{O}\times\mathcal{X}_{U}\times\mathcal{N}_{O}\times\mathcal{N}_{U}\rightarrow\mathcal{N}_{O}\times\mathcal{N}_{U}$.
Having observed a single draw $\left(X_{O},N_{O}\right)$ from the distribution $\mathcal{P}_{O}$, we are then interested in what happens to the endogenous variables $\left(N_{O},N_{U}\right)$ when we change the baseline exogenous variables $\left(X_{O},X_{U}\right)$ in a proportional way. Formally, for a given vector of exogenous change variables $\left(\hat{X}_{O},\hat{X}_{U}\right)$, we want to find a corresponding vector of endogenous change variables $\left(\hat{N}_{O},\hat{N}_{U}\right)$ such that
$f\left(X_{O}\odot\hat{X}_{O},X_{U}\odot\hat{X}_{U},N_{O}\odot\hat{N}_{O},N_{U}\odot\hat{N}_{U};\theta\right)=\boldsymbol{0},$
where $\odot$ denotes element-wise multiplication. The ultimate object of interest, which I denote by $\hat{k}$, will then be some transformation of the endogenous change variables $\left(\hat{N}_{O},\hat{N}_{U}\right)$, the observables $\left(X_{O},N_{O}\right)$ and the structural parameter $\theta$. It could for example be a specific endogenous change variable or a change in welfare. The key assumption that this counterfactual change variable of interest has to satisfy is:
Assumption 1.
For a given counterfactual question of interest, we can write $\hat{k}$ as a function solely of the observables $\left(X_{O},N_{O}\right)$:
$\hat{k}$ $=g_{\hat{k}}\left(X_{O},N_{O};\theta\right),$ (2)
for some known function $g_{\hat{k}}:\mathcal{X}_{O}\times\mathcal{N}_{O}\rightarrow\mathbb{R}$.
Assumption 1 implies that if $\left(X_{O},N_{O}\right)$ is observed without error and the structural parameter $\theta$ is known, we can perfectly recover $\hat{k}$. Loosely, in quantitative spatial models Assumption 1 often holds because, for a fixed counterfactual question of interest as described by $\left(\hat{X}_{O},\hat{X}_{U}\right)$, we can solve for the endogenous change variables $\left(\hat{N}_{O},\hat{N}_{U}\right)$ as a function of the observables $\left(X_{O},N_{O}\right)$. The exact functional form of $g_{\hat{k}}$ depends on the specific quantitative spatial model that is considered. In Appendix A I discuss the relevant underlying assumptions for two leading classes of models, namely invertible models and exact hat algebra models.
2.1.1 Running Example: Armington Model
For illustration, I will now work through finding $g_{\hat{k}}$ in a simple example. The simplest workhorse model in trade economics is the Armington model (Armington, 1969), as for example outlined in Costinot and Rodríguez-Clare (2014). There are countries indexed by $i,j=1,...,n$, which make a distinct good with productivity $Q_{i}$. Preferences are CES, so utility, which equals real consumption, for country $j$ is given by
$C_{j}=\left(\sum_{i=1}^{n}C_{ij}^{\frac{\varepsilon}{\varepsilon+1}}\right)^{\frac{\varepsilon+1}{\varepsilon}},\qquad j=1,...,n.$
Here, $C_{ij}$ is the demand for good $i$ in country $j$ and $\varepsilon>0$ is the trade elasticity, which equals the elasticity of substitution between goods from different countries minus one. Trade between countries is subject to iceberg trade costs, which means that in order to sell one unit of a good in country $j$, country $i$ must ship $\tau_{ij}\geq 1$ units, with $\tau_{ii}=1$. Assuming perfect competition, it follows that the price of the good from country $i$ in country $j$ is $P_{ij}=Y_{i}\tau_{ij}/Q_{i}$, where $Y_{i}$ is the total income in country $i$.
The relevant exogenous variables are then endowments $\left\{Q_{i}\right\}$ and trade costs $\left\{\tau_{ij}\right\}$. The endogenous variables are income levels $\left\{Y_{i}\right\}$ and expenditure shares $\left\{\lambda_{ij}\right\}$. Hence, using the notation outlined above, we have
[…]
$X_{U}$ =($=\left(\left\{Q_{i}\right\},\left\{\tau_{ij}\right\}\right)$
$N_{O}$ =($=\left(\left\{Y_{i}\right\},\left\{\lambda_{ij}\right\}\right)$
[…]
$\theta$ $=\varepsilon.$
It can be shown that the equilibrium conditions are:
$Y_{i}$ $=\sum_{j}\lambda_{ij}Y_{j},\qquad i=1,...,n,$ (3)
$\lambda_{ij}$ $=\frac{\left(\tau_{ij}Y_{i}\right)^{-\varepsilon}Q_{i}^{\varepsilon}}{\sum_{k}\left(\tau_{kj}Y_{k}\right)^{-\varepsilon}Q_{k}^{\varepsilon}},\qquad i,j=1,...,n.$ (4)
These equations can be rearranged to find the function $f$ as in Equation (1). We can plug the second set of equations into the first set of equations and solve for income levels $\left\{Y_{i}\right\}$ as functions of the exogenous variables. By Walras’ Law, income levels are only pinned down up to a multiplicative constant. Having solved for $\left\{Y_{i}\right\}$, we can then also solve for $\left\{\lambda_{ij}\right\}$. The model exactly matches the data in that for given income levels $\left\{Y_{i}\right\}$ and expenditure shares $\left\{\lambda_{ij}\right\}$, there exist trade costs $\left\{\tau_{ij}\right\}$ and productivities $\left\{Q_{i}\right\}$ such that these equilibrium conditions hold exactly.
Now, say we are interested in the counterfactual where we change the trade costs $\left\{\tau_{ij}\right\}$ proportionally by $\left\{\hat{\tau}_{ij}\right\}$, holding productivities $\left\{Q_{i}\right\}$ constant. The unobserved counterfactual levels of the trade costs hence are $\left\{\tau_{ij}\hat{\tau}_{ij}\right\}$. In Appendix B.1 it is shown that we can find:
$\hat{Y}_{i}Y_{i}$ $=\sum_{j}\hat{\lambda}_{ij}\lambda_{ij}\hat{Y}_{j}Y_{j},\qquad i=1,...,n,$ (5)
$\hat{\lambda}_{ij}$ $=\frac{\left(\hat{\tau}_{ij}\hat{Y}_{i}\right)^{-\varepsilon}}{\sum_{k}\left(\hat{\tau}_{kj}\hat{Y}_{k}\right)^{-\varepsilon}\lambda_{kj}},\qquad i,j=1,...,n.$
By again plugging in the second set of equations into the first one, we can then solve for the changes in income levels $\left\{\hat{Y}_{i}\right\}$ which again, by Walras’ Law, are only pinned down up to a multiplicative constant, and subsequently solve for $\left\{\hat{\lambda}_{ij}\right\}$. Hence, for a fixed counterfactual question of interest $\left\{\hat{\tau}_{ij}\right\}$, we can solve for $\hat{N}_{O}=\left(\left\{\hat{Y}_{i}\right\},\left\{\hat{\lambda}_{ij}\right\}\right)$ as a function solely of the observables $N_{O}=\left(\left\{Y_{i}\right\},\left\{\lambda_{ij}\right\}\right)$ and the structural parameter $\theta=\varepsilon$.
The counterfactual change variables of interest are the changes in welfare, which in this case correspond to the changes in real consumption $\left\{\hat{C}_{i}\right\}$. In Costinot and Rodríguez-Clare (2014), it is shown that these changes in welfare are
$\hat{C}_{i}=\hat{\lambda}_{ii}^{-1/\varepsilon},\qquad i=1,...,n.$
It follows that for each of these counterfactual change variables of interest, for a given counterfactual question as described by $\left\{\hat{\tau}_{ij}\right\}$, we only require knowledge of the observables $\left\{Y_{i}\right\}$ and $\left\{\lambda_{ij}\right\}$. For example, focusing on the change in welfare in the first country, if we assume that the trade elasticity $\varepsilon$ is known we find
$\hat{k}=\hat{C}_{1}=g_{\hat{C}_{1}}\left(\left\{\lambda_{ij}\right\},\left\{Y_{i}\right\};\varepsilon\right).$ (6)
2.2 Incorporating Parameter Uncertainty
The counterfactual change variable of interest will generally depend on the structural parameter $\theta$. If we assume that this structural parameter is fixed and known, our object of interest is $\hat{k}=g_{\hat{k}}\left(X_{O},N_{O};\theta\right)$. In practice, however, we do not know the structural parameter exactly and we hence have to use an estimator $\tilde{\theta}$. This subsection shows that the notation from Assumption 1 is general, and can account for estimation error by finding the appropriate $g$-functions.
Suppose that, instead of the true parameter $\theta$, we use an estimator $\tilde{\theta}\left(X_{O},N_{O}\right)$. Our focus then shifts from the true counterfactual change variable of interest $\hat{k}$ to the point estimate
$\hat{k}^{E}=g_{\hat{k}^{E}}\left(X_{O},N_{O}\right)\equiv g_{\hat{k}}\left(X_{O},N_{O};\tilde{\theta}\left(X_{O},N_{O}\right)\right).$
If we know the sampling distribution of $\tilde{\theta}\left(X_{O},N_{O}\right)$, we can find a $100\left(1-\beta\right)\%$ frequentist confidence region around $\hat{k}^{E}$, which I will denote by $\left[L_{\hat{k}^{E}}^{\beta},U_{\hat{k}^{E}}^{\beta}\right]$. Our object of interest then becomes this confidence region, and it follows that the relevant $g$-functions are
$L_{\hat{k}^{E}}^{\beta}$ $=g_{L_{\hat{k}^{E}}^{\beta}}\left(X_{O},N_{O}\right)\quad\mathrm{and}\quad U_{\hat{k}^{E}}^{\beta}=g_{U_{\hat{k}^{E}}^{\beta}}\left(X_{O},N_{O}\right).$
In practice it is common to assume there is no uncertainty around estimators for the structural parameter $\theta$, both in the case where estimators are taken from other papers and in the case where estimators depend on the data. An exception is Adao, Costinot, and Donaldson (2017), which reports $\left[L_{\hat{k}^{E}}^{\beta},U_{\hat{k}^{E}}^{\beta}\right]$ for the counterfactual change variables of interest.
A leading case is where the sampling distribution is approximately normal with variance-covariance matrix $\tilde{\Sigma}\left(\tilde{\theta}\left(X_{O},N_{O}\right)\right)$ and $g_{\hat{k}}$ is smooth in $\theta$ so that we can use the delta method. [^2] This notation nests the scenario where we use an estimator from another study that used different data. In that case $\tilde{\theta}$ is independent from $\left(X_{O},N_{O}\right)$ and we would write $\tilde{\Sigma}\left(\tilde{\theta}\right)$. If $g_{\hat{k}}$ is not smooth in $\theta$ one can use the bootstrap by obtaining draws $\tilde{\theta}_{b}\sim\mathcal{N}\left(\tilde{\theta},\tilde{\Sigma}\left(\tilde{\theta}\left(X_{O},N_{O}\right)\right)\right)$ and calculating $\tilde{k}_{b}^{E}=g_{\hat{k}}\left(X_{O},N_{O};\tilde{\theta}_{b}\right)$ for $b=1,...,B$. We can then obtain a $100\left(1-\beta\right)\%$ frequentist confidence region by finding the $\beta/2$ and $21-\beta/$ 2 quantiles of $\left\{\tilde{k}_{b}^{E}\right\}$. This yields
$\tilde{k}^{E}\sim\mathcal{N}\left(\hat{k}^{E},\frac{\partial g_{\hat{k}}}{\delta\theta}|_{\theta=\tilde{\theta}\left(X_{O},N_{O}\right)}\tilde{\Sigma}\left(\tilde{\theta}\left(X_{O},N_{O}\right)\right)\frac{\partial g_{\hat{k}}}{\delta\theta}|_{\theta=\tilde{\theta}\left(X_{O},N_{O}\right)}^{{}^{\prime}}\right)\equiv\mathcal{N}\left(\hat{k}^{E},\tilde{\Sigma}\left(\tilde{k}^{E}\right)\right),$
from which we can obtain the $100\left(1-\beta\right)\%$ frequentist confidence region
$\left[L_{\hat{k}^{E}}^{\beta},U_{\hat{k}^{E}}^{\beta}\right]=\left[\tilde{k}^{E}-z_{1-\beta/2}\sqrt{\tilde{\Sigma}\left(\tilde{k}^{E}\right)},\tilde{k}^{E}+z_{1-\beta/2}\sqrt{\tilde{\Sigma}\left(\tilde{k}^{E}\right)}\right].$
2.2.1 Running Example: Armington Model (Continued)
For the Armington model, trade flows $\left\{F_{ij}\right\}$ are the underlying data that determine the expenditure shares through $\lambda_{ij}=\frac{F_{ij}}{\sum_{\ell}F_{\ell j}}$. These trade flows are also used to estimate the trade elasticity $\varepsilon$. It follows that our object of interest is
$\hat{C}_{1}^{E}=g_{\hat{C}_{1}^{E}}\left(\left\{F_{ij}\right\},\left\{Y_{i}\right\}\right)\equiv g_{\hat{C}_{1}}\left(\left\{F_{ij}\right\},\left\{Y_{i}\right\};\tilde{\varepsilon}\left(\left\{F_{ij}\right\}\right)\right).$
Here, as for example outlined in Allen, Arkolakis, and Takahashi (2020), $\tilde{\varepsilon}\left(\left\{F_{ij}\right\}\right)$ is an estimator obtained from
$\log\frac{F_{ij}}{F_{jj}}=-\varepsilon\log\tilde{\tau}_{ij}+\gamma_{i}+\gamma_{j}+\phi_{ij},$ (7)
for $\tilde{\tau}_{ij}$ a proxy for the trade costs between country $i$ and $j$, $\gamma_{i}$ the country fixed effect of country $i$ and $\phi_{ij}$ an error term. [^3] An example for such a proxy is given in Allen, Arkolakis, and Takahashi (2020), which assumes that $-\varepsilon\log\tau_{ij}$ can be written as a function of their continents of origin and destination and the decile of distance between countries $i$ and $j$. From this regression, we can also learn $\tilde{\Sigma}\left(\tilde{\varepsilon}\left(\left\{F_{ij}\right\}\right)\right)$ and in turn find $\left[L_{\hat{C}_{1}^{E}}^{\beta},U_{\hat{C}_{1}^{E}}^{\beta}\right]$. It follows that our relevant $g$-functions are
$L_{\hat{C}_{1}^{E}}^{\beta}$ $=g_{L_{\hat{C}_{1}^{E}}^{\beta}}\left(\left\{F_{ij}\right\},\left\{Y_{i}\right\}\right)\quad\mathrm{and}\quad U_{\hat{C}_{1}^{E}}^{\beta}=g_{U_{\hat{C}_{1}^{E}}^{\beta}}\left(\left\{F_{ij}\right\},\left\{Y_{i}\right\}\right).$ (8)
3 Empirical Bayes Uncertainty Quantification
This section introduces measurement error to quantitative spatial models and discusses the differences with the existing literature on measurement error. It introduces an empirical Bayes approach to uncertainty quantification in the presence of measurement error, and outlines how to obtain credible sets for the counterfactual change variable of interest.
3.1 Introducing Measurement Error
Section 2 argued that our object of interest could be either $\hat{k}$, $\hat{k}^{E}$, $L_{\hat{k}^{E}}^{\beta}$ or $U_{\hat{k}^{E}}^{\beta}$. Furthermore, under Assumption 1, we saw that we can write the object of interest as a function solely of the observables $\left(X_{O},N_{O}\right)$. Hence, going forward, I will use the notation
$T=g_{T}\left(X_{O},N_{O}\right),$ (9)
where $g_{T}$ is a known function and, depending on the context, $T$ can equal $\hat{k}$, $\hat{k}^{E}$, $L_{\hat{k}^{E}}^{\beta}$ or $U_{\hat{k}^{E}}^{\beta}$.
The fact that $T$ only depends on the observables is convenient for answering counterfactual questions. However, the observables are economic variables which are often measured with error. For instance, Ortiz-Ospina and Beltekian (2018) and Goes (2023) highlight that there are large discrepancies between and within various data sources from trade and international economics. Motivated by this I assume that, instead of the true $X_{O}$ and $N_{O}$, we observe noisy versions $\tilde{X}_{O}$ and $\tilde{N}_{O}$, respectively, where
$\left(\tilde{X}_{O},\tilde{N}_{O}\right)|\left(X_{O},N_{O}\right)=m\left(X_{O},N_{O},\xi\right),$ (10)
for some independent measurement error variable $\xi\in\Xi\subseteq\mathbb{R}^{d_{\xi}}$ and measurement error function $m:\mathcal{X}_{O}\times\mathcal{N}_{O}\times\Xi\rightarrow\mathcal{X}_{O}\times\mathcal{N}_{O}$. I will assume that $m\left(\cdot\right)$ is known and will treat the distribution of $\xi$ as something to be learned.
Since the object of interest $T$ is a function of the true data, we will have uncertainty around $T$ when the data are noisily measured:
$\tilde{T}=g_{T}\left(\tilde{X}_{O},\tilde{N}_{O}\right).$
Our goal is to quantify uncertainty about $T$, and in particular to provide sets that cover $T$ with prescribed probability.
3.1.1 Relation to Measurement Error Literature
The literature on measurement error in nonlinear models is extensive, as reviewed in Hu (2015) and Schennach (2016), and the most closely related strand of measurement error literature is that on nonseparable error models (Matzkin, 2003; Chesher, 2003; Hoderlein and Mammen, 2007; Matzkin, 2008; Hu and Schennach, 2008; Schennach, White, and Chalak, 2012; Song, Schennach, and White, 2015). However, conventional results do not apply to my setting. The key distinguishing feature of the setting in this paper is that the object of interest $T=g_{T}\left(X_{O},N_{O};\theta\right)$ directly depends on the correctly measured data, because the equality in Assumption 1 is an exact statement. In contrast, in conventional measurement error settings the object of interest is a function of the correctly measured distribution of the data, $\mathcal{P}_{O}$, rather than the actual observations, $\left(X_{O},N_{O}\right)$. This distinction is important because in my setting, even if we could perfectly estimate $\mathcal{P}_{O}$, measurement error would still be a problem. For example in the Armington model, to answer counterfactual questions we need the realized trade flows, rather than the distribution of trade flows. By virtue of that, we need to account for uncertainty about the underlying observations themselves, which is a different problem than that in the conventional measurement error literature. This reliance on the individual observations rather than the distribution is a key distinction, and it is caused by the underlying assumption that the model exactly matches the data.
3.1.2 Running Example: Armington Model (Continued)
For the Armington model, I will assume that there is measurement error in bilateral trade flows $\left\{F_{ij}\right\}$. This is plausible because in Linsi, Burgoon, and Mügge (2023) it is shown that there are so-called mirror discrepancies in bilateral trade flows between almost all countries. This means that, for instance, while the value that Germany reports it imported from France and the value that France reports it exported to Germany should be the same, in practice they are often different. Instead of the true trade flows we observe noisy trade flows $\left\{\tilde{F}_{ij}\right\}$ which then lead to noisy expenditure shares $\left\{\tilde{\lambda}_{ij}\right\}$. Since trade flows are non-negative, I assume measurement error is log-normally distributed:
$\log\tilde{F}_{ij}|\log F_{ij}\sim\mathcal{N}\left(\log F_{ij},\varsigma^{2}\right),\qquad i,j=1,...,n.$
In the case where we also consider estimation error, this means that instead of $L_{\hat{C}_{1}^{E}}^{\beta}$ and $U_{\hat{C}_{1}^{E}}^{\beta}$ we observe
$\tilde{L}_{\hat{C}_{1}^{E}}^{\beta}$ $=g_{L_{\hat{C}_{1}^{E}}^{\beta}}\left(\left\{\tilde{F}_{ij}\right\},\left\{Y_{i}\right\}\right)\quad\mathrm{and}\quad\tilde{U}_{\hat{C}_{1}^{E}}^{\beta}=g_{U_{\hat{C}_{1}^{E}}^{\beta}}\left(\left\{\tilde{F}_{ij}\right\},\left\{Y_{i}\right\}\right).$ (11)
3.2 Motivation for Empirical Bayes Approach
The previous section introduced measurement error to quantitative spatial models. We are interested in $T=g_{T}\left(X_{O},N_{O}\right)$ but instead observe $\tilde{T}=g_{T}\left(\tilde{X}_{O},\tilde{N}_{O}\right)$, which leads to the goal of generating an interval which covers $T$ with prescribed probability.
To achieve this goal, I take an empirical Bayes (EB) approach, and introduce a model for the measurement error and estimate a prior for the true underlying data. [^4] Rather than estimating the parameters of the prior for the true underlying data, which corresponds to an EB approach, one could alternatively specify priors for these parameters, which corresponds to a hierarchical Bayes approach. Using Bayes’ rule, the measurement error model and the prior can be combined to find the posterior of the true data given the noisy data. We can then feed this posterior through $T=g_{T}\left(X_{O},N_{O}\right)$ to obtain the posterior of the counterfactual change variable of interest.
A first reason for using an EB approach is that it allows researchers to incorporate economic knowledge through the prior. For example when considering measurement error in positive flows between locations one can fit a gravity model as a prior, which I will do in Section 3.4. A second reason is that, given the prior, the resulting procedure is automated and computationally appealing.
An alternative approach, which I call an "adversarial approach" following the machine learning literature, and which is developed in Appendix C, prespecifies a bound on the measurement error and generates worst-case bounds on the counterfactual. This avoids the need to fully specify a prior distribution for the measurement error, at the cost of having to prespecify a bound on the measurement error. The adversarial approach is more computationally demanding and, to be feasible in high-dimensional applications, must rely on approximations.
3.3 Obtaining Credible Sets for $T$
I now discuss how to obtain credible sets for $T=g_{T}\left(X_{O},N_{O}\right)$ when we observe $\left(\tilde{X}_{O},\tilde{N}_{O}\right)$. That is, denoting with $\mathbb{P}_{\pi}$ the probability under the prior, for a given $\gamma\in\left[0,1\right]$ we want to find a region $\mathcal{C}_{T}^{\gamma}$ such that $\mathbb{P}_{\pi}\left[T\in\mathcal{C}_{T}^{\gamma}|\tilde{X}_{O},\tilde{N}_{O}\right]\geq 1-\gamma.$ Towards that end, given a prior on the true data and a model for the measurement error,
$\begin{cases}\begin{array}[]{c}\mathrm{prior:}\\ \mathrm{measurement\>error:}\end{array}&\begin{array}[]{c}\pi\left(X_{O},N_{O}\right)\\ \pi\left(\tilde{X}_{O},\tilde{N}_{O}|X_{O},N_{O}\right),\end{array}\end{cases}$
we can use Bayes’ rule to obtain the posterior of the true data given the noisy data
[…]
This posterior then allows us to generate draws from our posterior for the true data given the noisy data. For each of these draws, we can calculate the corresponding value for the counterfactual change variable of interest using the relationship $T=g_{T}\left(X_{O},N_{O}\right)$. This allows us to find the posterior of the true object of interest $T$ given the noisy data,
$\pi\left(T|\tilde{X}_{O},\tilde{N}_{O}\right).$
Once we have this posterior, we can obtain credible sets by for example taking an interval between two relevant quantiles. Note that these credible sets can not be interpreted as frequentist confidence intervals, and are dependent on the specified prior distribution.
Proposition 1.
Denote with $q_{\alpha}$ the $\alpha$-quantile of the posterior $T|\tilde{X}_{O},\tilde{N}_{O}$. A $100\left(1-\gamma\right)\%$ equal-tailed credible set is given by $\mathcal{C}_{T}^{\gamma}=\left[q_{\gamma/2},q_{1-\gamma/2}\right]$:
$\mathbb{P}_{\pi}\left[T\in\mathcal{C}_{T}^{\gamma}|\tilde{X}_{O},\tilde{N}_{O}\right]\geq 1-\gamma.$
Remark 1.
If we want to account for both estimation error and measurement error simultaneously, we focus on covering the $100\left(1-\beta\right)\%$ frequentist confidence region $\left[L_{\hat{k}^{E}}^{\beta},U_{\hat{k}^{E}}^{\beta}\right]$ with prescribed probability. We then first choose $T$ equal to $L_{\hat{k}^{E}}^{\beta}$ and then to $U_{\hat{k}^{E}}^{\beta}$. Denote with $\ell_{\alpha}^{\beta}$ and $\upsilon_{\alpha}^{\beta}$ the $\alpha$-quantiles of the resulting posteriors $L_{\hat{k}^{E}}^{\beta}|\tilde{X}_{O},\tilde{N}_{O}$ and $U_{\hat{k}^{E}}^{\beta}|\tilde{X}_{O},\tilde{N}_{O}$, respectively. In that case, a $100\left(1-\beta-\gamma\right)\%$ equal-tailed credible set is given by $\mathcal{C}_{\hat{k}^{E}}^{\beta,\gamma}=\left[\ell_{\gamma/2}^{\beta},\upsilon_{1-\gamma/2}^{\beta}\right]$.
3.4 Widely Applicable Default Approach
To provide a widely applicable default approach, consider the setting where we can write $T=g_{T}\left(\left\{F_{ij}\right\}\right)$ for a known function $g_{T}$ and a set of positive flows between locations, $\left\{F_{ij}\right\}$, a setup that is commonplace in quantitative spatial models (Costinot and Rodríguez-Clare, 2014; Redding and Rossi-Hansberg, 2017). Given the observed noisy flows $\left\{\tilde{F}_{ij}\right\}$, our goal is to find the posterior distribution $T|\left\{\tilde{F}_{ij}\right\}$. This requires a measurement error model and a prior on the true flows $\left\{F_{ij}\right\}$.
In the default approach I recommend, both the prior on the true flows and the measurement errors are assumed to be log-normally distributed. This implies that posterior of the true flows given the noisy flows will also be log-normally distributed. This conjugacy is convenient, as it is computationally cheap to draw from the normal distribution. [^5] One could potentially enrich this by using a mixture of normals, in which case we will still have conjugacy. Furthermore, I assume that the prior mean exhibits a gravity relationship, for which there is strong empirical evidence, as for example outlined in Allen and Arkolakis (2018). This is summarized in the following assumption:
Assumption 2.
We have
$\begin{cases}\begin{array}[]{c}\mathrm{prior:}\\ \mathrm{measurement\>error:}\end{array}&\begin{array}[]{c}\log F_{ij}\sim\mathcal{N}\left(\beta\log\mathrm{dist}_{ij}+\alpha_{i}^{\mathrm{orig}}+\alpha_{j}^{\mathrm{dest}},s^{2}\right)\\ \log\tilde{F}_{ij}|\log F_{ij}\sim\mathcal{N}\left(\log F_{ij},\varsigma^{2}\right),\end{array}\end{cases}$ (12)
for $i,j=1,...,n$, where $\mathrm{dist}_{ij}$ denotes the distance between locations $i$ and $j$, $\alpha_{i}^{\mathrm{orig}}$ is an origin fixed effect and $\alpha_{j}^{\mathrm{dest}}$ is a destination fixed effect.
It follows that the posterior for the true log flow between location $i$ and $j$, $\log F_{ij}$, given its noisy version, $\log\tilde{F}_{ij}$ is given by
$\mathcal{N}\left(\frac{s^{2}}{s^{2}+\varsigma^{2}}\log\tilde{F}_{ij}+\frac{\varsigma^{2}}{s^{2}+\varsigma^{2}}\left\{\beta\log\mathrm{dist}+\alpha_{i}^{\mathrm{orig}}+\alpha_{j}^{\mathrm{dest}}\right\},\left(\frac{1}{s^{2}}+\frac{1}{\varsigma^{2}}\right)^{-1}\right),$
for $i,j=1,...,n$. Then, a default procedure for obtaining a credible set for $T$ is summarized in the following algorithm:
Algorithm 1 Credible set for $T=g_{T}\left(\left\{F_{ij}\right\}\right)$
  1. 1.
    Estimate the parameters $\vartheta=\left(\beta,\left\{\alpha_{i}^{\mathrm{orig}}\right\},\left\{\alpha_{i}^{\mathrm{dest}}\right\},s^{2},\varsigma^{2}\right)$:
    1. (a)
      Find estimates $\left\{\hat{\varsigma}^{2}\right\}$ using a specific data structure (as in Section 4) or domain knowledge (as in Section 5).
    2. (b)
      Obtain estimates $\left(\hat{\beta},\left\{\hat{\alpha}_{i}^{\mathrm{orig}}\right\},\left\{\hat{\alpha}_{i}^{\mathrm{dest}}\right\}\right)$ from the regression $\log\tilde{F}_{ij}=\beta\log\mathrm{dist}_{ij}+\alpha_{i}^{\mathrm{orig}}+\alpha_{j}^{\mathrm{dest}}+\phi_{ij}$, for $\phi_{ij}$ an error term.
    3. (c)
      Find $\hat{s}^{2}=\hat{r}^{2}-\hat{\varsigma}^{2}$, for $\hat{r}^{2}$ the sample variance of the residuals $\left\{\log\tilde{F}_{ij}-\hat{\beta}\log\mathrm{dist}_{ij}-\hat{\alpha}_{i}^{\mathrm{orig}}-\hat{\alpha}_{j}^{\mathrm{dest}}\right\}_{i,j=1}^{n}$.
  2. 2.
    Take $B$ draws (choose $B$ and $\gamma$ such that $\gamma/2\cdot B\in\mathbb{N}$) from the estimated posterior $\log F_{ij}|\log\tilde{F}_{ij}$ that uses $\hat{\vartheta}$ for $i,j=1,...,n$ and indicate them by $\log F_{ij,1},...,\log F_{ij,B}$ for $i,j=1,...,n$.
  3. 3.
    For $b=1,...,B$, compute $T_{b}=g_{T}\left(\left\{F_{ij,b}\right\}_{i,j=1}^{n}\right)$.
  4. 4.
    Sort these draws to obtain $\left\{T^{(b)}\right\}_{i=1}^{B}$ with $T^{(1)}\leq T^{(2)}\leq...\leq T^{(B)}$.
  5. 5.
    Report
    $\mathcal{C}_{T}^{\gamma}=\left[T^{(\gamma/2\cdot B)},T^{(\left(1-\gamma/2\right)\cdot B)}\right].$
Remark 2.
When we observe a panel of flows $\left\{F_{ij,t}\right\}$ as in Section 4, we can estimate the measurement error variances and prior variances for each flow separately. The assumptions change to
$\begin{cases}\begin{array}[]{c}\mathrm{prior:}\\ \mathrm{measurement\>error:}\end{array}&\begin{array}[]{c}\log F_{ij,t}\sim\mathcal{N}\left(\beta_{t}\log\mathrm{dist}_{ij}+\alpha_{i,t}^{\mathrm{orig}}+\alpha_{j,t}^{\mathrm{dest}},s_{ij}^{2}\right)\\ \log\tilde{F}_{ij,t}|\log F_{ij,t}\sim\mathcal{N}\left(\log F_{ij,t},\varsigma_{ij}^{2}\right),\end{array}\end{cases}$ (13)
for $i,j=1,...,n$ and $t=1,...,\bar{t}$, which yields the posterior
$\mathcal{N}\left(\frac{s_{ij}^{2}}{s_{ij}^{2}+\varsigma_{ij}^{2}}\log\tilde{F}_{ij}+\frac{\varsigma_{ij}^{2}}{s_{ij}^{2}+\varsigma_{ij}^{2}}\left\{\beta_{t}\log\mathrm{dist}+\alpha_{i,t}^{\mathrm{orig}}+\alpha_{j,t}^{\mathrm{dest}}\right\},\left(\frac{1}{s_{ij}^{2}}+\frac{1}{\varsigma_{ij}^{2}}\right)^{-1}\right),$
for $i,j=1,...,n$. This implies that now $\vartheta=\left(\left\{\beta_{t}\right\},\left\{\alpha_{i,t}^{\mathrm{orig}}\right\},\left\{\alpha_{i,t}^{\mathrm{dest}}\right\},\left\{s_{ij}^{2}\right\},\left\{\varsigma_{ij}^{2}\right\}\right)$. For $t=1,...,\bar{t}$, the estimates $\left(\hat{\beta}_{t},\left\{\hat{\alpha}_{i,t}^{\mathrm{orig}}\right\},\left\{\hat{\alpha}_{i,t}^{\mathrm{dest}}\right\}\right)$ are obtained from the regression $\log\tilde{F}_{ij,t}=\beta_{t}\log\mathrm{dist}_{ij,t}+\alpha_{i,t}^{\mathrm{orig}}+\alpha_{j,t}^{\mathrm{dest}}+\phi_{ij,t}$, and we have $\hat{s}_{ij}^{2}=\hat{r}_{ij}^{2}-\hat{\varsigma}_{ij}^{2}$, where $\hat{r}_{ij}^{2}$ denotes the sample variance of the residuals $\left\{\log\tilde{F}_{ij,t}-\hat{\beta}_{t}\log\mathrm{dist}_{ij}-\hat{\alpha}_{i,t}^{\mathrm{orig}}-\hat{\alpha}_{j,t}^{\mathrm{dest}}\right\}_{t=1}^{\bar{t}}$.
Remark 3.
One can verify how reasonable the normality assumption on the prior is by comparing the histogram of the normalized residuals
$\left\{\frac{\log\tilde{F}_{ij}-\left\{\hat{\beta}\log\mathrm{dist}_{ij}+\hat{\alpha}_{i}^{\mathrm{orig}}+\hat{\alpha}_{j}^{\mathrm{dest}}\right\}}{\hat{s}}\right\}$
with the probability density function of a standard normal. To check the reasonableness of the gravity prior, we can look at the adjusted R-squared of the gravity model and, following Allen and Arkolakis (2018), plot the log flows against the log distance, after partitioning out the origin and destination fixed effects. In Appendices D and E I perform both these checks for my applications.
Remark 4.
Specifying a measurement error model is difficult and one might be worried about misspecification. For the normal-normal model, we can use results on prior density-ratio class robustness developed in Geweke and Petrella (1998) to find worst-case bounds over a neighborhood that contains distributions that are not too far away from the assumed normal distribution for the measurement error model. The details can be found in Appendix F.
3.4.1 Running Example: Armington Model (Continued)
Consider again the Armington model where trade flows are noisily measured. This fits the setting from Section 3.4 and the prior and measurement error model are as in Equation (12). We can apply Algorithm 1 in which the estimated posterior of the true trade flow $\log F_{ij}$ given the noisy trade flow $\log\tilde{F}_{ij}$ is
$\mathcal{N}\left(\frac{\hat{s}^{2}}{\hat{s}^{2}+\hat{\varsigma}^{2}}\log\tilde{F}_{ij}+\frac{\hat{\varsigma}^{2}}{\hat{s}^{2}+\hat{\varsigma}^{2}}\left\{\hat{\beta}\log\mathrm{dist}_{ij}+\hat{\alpha}_{i}^{\mathrm{orig}}+\hat{\alpha}_{j}^{\mathrm{dest}}\right\},\left(\frac{1}{\hat{s}^{2}}+\frac{1}{\hat{\varsigma}^{2}}\right)^{-1}\right),$ (14)
for $i,j=1,...,n$. Focusing again on the change in welfare of the first country which uses the estimated trade elasticity, $\hat{C}_{1}^{E}$, we can use the posterior in Equation (14) to find a credible set that covers $\left[L_{\hat{C}_{1}^{E}}^{\beta},U_{\hat{C}_{1}^{E}}^{\beta}\right]$ from Equation (8) with prescribed probability. Because we know that in turn $\hat{C}_{1}^{E}$ lies in $\left[L_{\hat{C}_{1}^{E}}^{\beta},U_{\hat{C}_{1}^{E}}^{\beta}\right]$ with prescribed probability, this allows us to then find a credible set for $\hat{C}_{1}^{E}$ as in Remark 1.
4 Application 1: Adao, Costinot, and Donaldson (2017)
4.1 Model and Counterfactual Question of Interest
The empirical application of Adao, Costinot, and Donaldson (2017) investigates the effects of China joining the WTO, the so-called China shock. Specifically, the authors examine what would have happened to China’s welfare if China’s trade costs would stay constant at their 1995 levels. They consider $n$ countries and $\bar{t}$ time periods. Denoting with $\tau_{ij,t}$ the trade cost between country $i$ and country $j$ in period $t$, this implies the following counterfactual question of interest:
$\hat{\tau}_{ij,t}$ $=\frac{\tau_{ij,95}}{\tau_{ij,t}},\quad\mathrm{if}\>i\>\mathrm{or}\>j\>\mathrm{is}\>\mathrm{China,}$
$\hat{\tau}_{ij,t}$ $=1,\quad\mathrm{otherwise.}$
Here, welfare is defined as the percentage change in income that the representative agent in China would be indifferent about accepting instead of the counterfactual change in trade costs from $\left\{\tau_{ij,t}\right\}$ to $\left\{\hat{\tau}_{ij,t}\tau_{ij,t}\right\}$. The details of the model can be found in Appendix D.1. The key insight is that we can express the estimated change in China’s welfare in period $t$, denoted by $\hat{W}_{\mathrm{China},t}^{E}$, as a function of all the bilateral trade flows in different periods $\left\{F_{ij,t}\right\}$ and the estimated trade elasticity. Since the estimated trade elasticity is itself a function of the bilateral trade flows, we can write
$\hat{W}_{\mathrm{China},t}^{E}=g_{\hat{W}_{\mathrm{China},t}}\left(\left\{F_{ij,t}\right\};\tilde{\varepsilon}\left(\left\{F_{ij,t}\right\}\right)\right),$ (15)
for known functions $\tilde{\varepsilon}:\mathbb{R}_{+}^{Tn\left(n-1\right)}\rightarrow\mathbb{R}$ and $g_{\hat{W}_{\mathrm{China},t}}:\mathbb{R}_{+}^{Tn\left(n-1\right)}\rightarrow\mathbb{R}$. We can then apply Algorithm 1 and Remark 1 to find a credible set for $\hat{W}_{\mathrm{China},t}^{E}$.
4.2 Measurement Error Model and Prior
The model fits the panel setting from Section 3.4 as outlined in Remark 2, and the prior and measurement error model are as in Equation (13). For estimation of $\vartheta$, I use the distance dataset $\left\{\mathrm{dist}_{ij}\right\}_{i\neq j}$ from Mayer and Zignago (2011) and the mirror trade dataset from Linsi, Burgoon, and Mügge (2023). This mirror trade dataset has two estimates of each bilateral trade flow, both as reported by the exporter and as by the importer. I interpret this as observing two independent noisy observations per time period for each bilateral trade flow: $\left\{\left\{\tilde{F}_{ij,t}^{1},\tilde{F}_{ij,t}^{2}\right\}_{t=1}^{T}\right\}_{i\neq j}$.
Then, to estimate $\varsigma_{ij}^{2}$, note that
$\log\tilde{F}_{ij,t}^{1}-\log\tilde{F}_{ij,t}^{2}\sim\mathcal{N}\left(0,2\varsigma_{ij}^{2}\right)$
for $i,j=1,...,n$, from which it follows that
$\hat{\varsigma}_{ij}^{2}=\frac{1}{2T}\sum_{t=1}^{T}\left(\log\tilde{F}_{ij,t}^{1}-\log\tilde{F}_{ij,t}^{2}\right)^{2}$
is an unbiased estimator of $\varsigma_{ij}^{2}$ for $i,j=1,...,n$. We can then estimate the rest of $\vartheta$ as outlined in Remark 2, adapting for the fact that we have two observations for each bilateral trade flow. For completeness, Appendix D.2 goes through the full procedure to obtain $\hat{\vartheta}=\left(\left\{\hat{\beta}_{t}\right\},\left\{\hat{\alpha}_{i,t}^{\mathrm{orig}}\right\},\left\{\hat{\alpha}_{i,t}^{\mathrm{dest}}\right\},\left\{\hat{s}_{ij}^{2}\right\},\left\{\hat{\varsigma}_{ij}^{2}\right\}\right)$.
To reduce the variability and leverage country information for $\left\{\hat{\varsigma}_{ij}^{2}\right\}$ and $\left\{\hat{s}_{ij}^{2}\right\}$, I fit the models:
$\hat{\varsigma}_{ij}^{2}=e^{\gamma_{i}^{\mathrm{exp}}+\gamma_{j}^{\mathrm{imp}}+u_{ij}^{\varsigma}}\quad\mathrm{and}\quad\hat{s}_{ij}^{2}=e^{\delta_{i}^{\mathrm{exp}}+\delta_{j}^{\mathrm{imp}}+u_{ij}^{s}},$ (16)
for $i,j=1,...,n$, with $\gamma_{i}^{\mathrm{exp}}$, $\gamma_{j}^{\mathrm{imp}}$, $\delta_{i}^{\mathrm{exp}}$ and $\delta_{i}^{\mathrm{imp}}$ country-exporter and country-importer fixed effects and $u_{ij}^{\varsigma}$ and $u_{ij}^{s}$ error terms. Then, rather than using $\hat{\varsigma}_{ij}^{2}$ and $\hat{s}_{ij}^{2}$ I will use the fitted values $\mathring{\varsigma}_{ij}^{2}=e^{\hat{\gamma}_{i}^{\mathrm{exp}}+\hat{\gamma}_{j}^{\mathrm{imp}}}$ and $\mathring{s}_{ij}^{2}=e^{\hat{\delta}_{i}^{\mathrm{exp}}+\hat{\delta}_{j}^{\mathrm{imp}}}$. Finally, the estimated posterior of the true trade flow $\log F_{ij,t}$ given the noisy trade flow $\log\tilde{F}_{ij,t}$ then is given, for $i,j=1,...,n$, by
$\mathcal{N}\left(\frac{\mathring{s}_{ij}^{2}}{\mathring{s}_{ij}^{2}+\mathring{\varsigma}_{ij}^{2}}\log\tilde{F}_{ij,t}+\frac{\mathring{\varsigma}_{ij}^{2}}{\mathring{s}_{ij}^{2}+\mathring{\varsigma}_{ij}^{2}}\left\{\hat{\beta}_{t}\mathrm{dist}_{ij}+\hat{\alpha}_{i,t}^{\mathrm{exp}}+\hat{\alpha}_{j,t}^{\mathrm{imp}}\right\},\left(\frac{1}{\mathring{s}_{ij}^{2}}+\frac{1}{\mathring{\varsigma}_{ij}^{2}}\right)^{-1}\right).$ (17)
4.3 Results
Having obtained a posterior for the true trade flows given the noisy trade flows, we can now find credible sets for the counterfactual change variables of interest. In Figure 1, I reproduce Figure 3 of Adao, Costinot, and Donaldson (2017), which plots the change in China’s welfare as a result of the China shock for each year in the period 1996-2011, and include 95% equal-tailed confidence intervals or credible sets. Specifically, I plot three different intervals.
The first is a frequentist confidence interval that only considers estimation error and hence assumes the data are perfectly measured. It is constructed using code provided by the authors, which uses a bootstrap procedure that samples from the asymptotic distribution of the GMM estimator for the trade elasticity $\varepsilon$. The resulting confidence intervals are small for the period before the year 2000, and then slowly become wider. For the period 2003-2011, with the exception of 2009, it becomes ambiguous whether joining the WTO had a positive or negative effect on China’s welfare.
The second region considers only measurement error and not sampling error in $\varepsilon$, so the goal is to find a 95% credible set for $\hat{W}_{\mathrm{China},t}^{E}$ for $t=1996,...,2011$. Towards this, I feed forward draws from the posterior in Equation (17) into the function $g_{\hat{W}_{\mathrm{China},t}}$ in Equation (15). The resulting credible set is slightly wider than the confidence interval based solely on estimation error, especially in the first few years. Again, for the period 2003-2011 it is unclear whether joining the WTO had a positive or negative effect on China’s welfare.
Finally, the third region combines estimation error and measurement error. As outlined in Remark 1, the focus then shifts to finding 97.5% credible sets for the upper and lower endpoints of a 97.5% frequentist confidence interval that considers only estimation error. The resulting lower bound seems to be a reasonable composition of the individual lower bounds considering only estimation error or only measurement error. On the other hand, the posterior distributions of the upper bounds have heavy right tails and the upper bounds become large.
Figure 1: EB uncertainty quantification for heteroskedastic normal shocks to $\left\{\log F_{ij,t}\right\}$ for the change in China’s welfare due to the China shock. The solid blue line is the estimate as reported in Adao, Costinot, and Donaldson (2017).
We can learn more from the posterior distributions than just credible sets. It might for example be of interest what the posterior probability is that the welfare in China increased as a result of the China shock, which is calculated in Appendix D.3. There, I also perform some additional analyses to check the robustness of these results.
5 Application 2: Allen and Arkolakis (2022)
5.1 Model and Counterfactual Question of Interest
The empirical application in Allen and Arkolakis (2022) aims to answer the question what the returns on investment are of all the highway segments of the US Interstate Highway network. The authors do so by introducing an economic geography model and calculating what happens to welfare after a small ($1\%$) improvement to all highway links. Combining these counterfactual welfare changes with how many lane-miles must be added in order to achieve the $1\%$ improvement, they find the highway segments with the greatest return on investment.
This exercise requires only data on incomes and traffic flows of the $n$ locations and knowledge of four structural model parameters. Three of these parameters are taken from the literature and assumed to have no uncertainty around them. The fourth, which is the strength of traffic congestion $\delta$, is estimated using the noisily measured traffic flow data. The details of the model can be found in Appendix E.1, but the key relation is the one that maps the average annual daily traffic (AADT) flows $\left\{F_{ij}\right\}$ to estimated welfare $\hat{W}^{E}$, which is
$\hat{W}^{E}=g_{\hat{W}}\left(\left\{F_{ij}\right\};\tilde{\delta}\left(\left\{F_{ij}\right\}\right)\right)$
for known functions $\tilde{\delta}:\mathbb{R}_{+}^{n\left(n-1\right)}\rightarrow\mathbb{R}$ and $g_{\hat{W}}:\mathbb{R}_{+}^{n\left(n-1\right)}\rightarrow\mathbb{R}$.
Similarly as for the application in Section 4, I distinguish between the scenarios where the strength of traffic congestion $\delta$ is estimated with and without estimation error. In Allen and Arkolakis (2022), estimation error is not considered so to obtain the frequentist confidence interval that accounts for estimation error I use the delta method as outlined in Section 2.2.
5.2 Measurement Error Model and Prior
Towards uncertainty quantification, I will add log-normal measurement error to non-zero traffic flows. Musunuru and Porter (2019) estimates that the measurement error variance of the logarithm of the average annual daily traffic (AADT) flows, which is exactly the data that Allen and Arkolakis (2022) uses, is between 0.05 and 0.20. To obtain a lower bound on uncertainty, I will use a uniform measurement error variance of $0.05$. I follow Musunuru and Porter (2019) by adding independent and identically distributed measurement error shocks to log traffic flows. Hence, this fits the setting from Section 3.4 and the prior and measurement error model are as in Equation (12).
Using $\hat{\varsigma}^{2}=0.05$, I estimate a prior variance of $\hat{s}^{2}=0.101$. This results in the following estimated posterior for the true traffic flow between country $i$ and $j$, $\log F_{ij}$, given its noisy version $\log\tilde{F}_{ij}$, for $i,j=1,...,n$,
$\mathcal{N}\left(0.669\cdot\log\tilde{F}_{ij}+0.331\cdot\left\{\hat{\beta}\mathrm{dist}_{ij}+\hat{\alpha}_{i}^{\mathrm{orig}}+\hat{\alpha}_{j}^{\mathrm{dest}}\right\},0.033\right).$
5.3 Results
The counterfactual question of interest is which links have the highest return on investment, and the authors of Allen and Arkolakis (2022) report the top ten links. For exposition, I will focus my analysis on the three best performing links. Similarly to the application in Section 4, I consider scenarios with only estimation error, only measurement error, and both estimation and measurement error. Table 1 shows the $95\%$ equal-tailed confidence intervals and credible sets for the top three links. Accounting only for estimation error results in bounds that are around $50\%$ away from the point estimate, and the bounds resulting from accounting only for measurement error are generally less wide.
Link 1 Link 2 Link 3
Point estimate 10.43 9.54 7.31
Only est. error [5.31, 15.55] [4.69, 14.38] [3.62, 11.00]
Only meas. error, EB [8.52, 14.05] [7.42, 10.96] [6.75, 8.18]
Est. and meas. error, EB [2.97, 24.14] [2.24 , 19.31] [1.99, 14.46]
Table 1: EB uncertainty quantification for the three links from Allen and Arkolakis (2022) with the highest return on investment. Link 1 is Kingsport-Bristol (TN-VA) to Johnson City (TN), link 2 is Greensboro-High Point (NC) to Winston-Salem (NC) and link 3 is Rochester (NY) to Batavia (NY).
From a policy perspective it is of interest whether the ranking between these links changes. Therefore, Table 2 shows the $95\%$ confidence intervals and credible sets for the difference between link 1 and link 2, and the difference between link 2 and link 3. When we consider estimation error, either separately or combined with measurement error, we fail to reject that these differences are zero and hence lose the ability to rank these three links. When we consider only measurement error, we can reject that these differences are zero. Additional discussion and analyses can be found in Appendices E.2 and E.3.
Link 1-Link 2 Link 2-Link 3
Point estimate 0.89 2.23
Only est. error [-6.16, 7.94] [-3.86, 8.32]
Only meas. error, EB [0.39, 5.25] [0.01, 3.31]
Est. and meas. error, EB [-408.36, 411.18] [-396.06, 397.33]
Table 2: EB uncertainty quantification for the differences between the three links from Allen and Arkolakis (2022) with the highest return on investment. Link 1 is Kingsport-Bristol (TN-VA) to Johnson City (TN), link 2 is Greensboro-High Point (NC) to Winston-Salem (NC) and link 3 is Rochester (NY) to Batavia (NY).
6 Conclusion
In this paper, I provide an econometric framework for examining the effect of parameter uncertainty and measurement error for an important class of quantitative spatial models. This setting differs from conventional measurement error models, because the object of interest directly depends on the correctly measured data rather than on the correctly measured distribution of the data. I take an empirical Bayes approach to uncertainty quantification and show how to obtain credible sets for the counterfactual change variables of interest. The proposed method accounts for the fact that the structural parameter is often estimated using the noisy data. I discuss the implications of considering measurement error and estimation error for applications in Adao, Costinot, and Donaldson (2017) and Allen and Arkolakis (2022). For both papers, I find substantial uncertainty in important economic quantities, which highlights the importance of uncertainty quantification.
Appendix A Finding $g_{\hat{k}}$ in Two Leading Classes of Models
This section discusses how to find the function $g_{\hat{k}}$ for two leading classes of models, and introduces two assumptions on the equilibrium conditions that commonly underly counterfactual analyses.
A.1 Assumptions
The two leading classes of quantitative spatial models I will consider are invertible models and exact hat algebra models. In these models, the following two assumptions on the equilibrium conditions $f$ hold:
Assumption 3 (Uniqueness).
For any $\left(X_{O},X_{U},\theta\right)$, there exists a (possibly up to a multiplicative constant) unique $\left(N_{O},N_{U}\right)$ such that (1) holds exactly. So for
$\mathcal{S}_{\left(f,X_{O},X_{U},\theta\right)}\equiv\left\{\left(N_{O},N_{U}\right)\in\mathcal{N}_{O}\times\mathcal{N}_{U}:f\left(X_{O},X_{U},N_{O},N_{U};\theta\right)=\boldsymbol{0}\right\},$
we have that $\left(N_{O}^{1},N_{U}^{1}\right),\left(N_{O}^{2},N_{U}^{2}\right)\in\mathcal{S}_{\left(f,X_{O},X_{U},\theta\right)}$ implies $\left(N_{O}^{1},N_{U}^{1}\right)=c\cdot\left(N_{O}^{2},N_{U}^{2}\right)$ for some $c>0$, for all $\left(X_{O},X_{U},\theta\right)\in\mathcal{X}_{O}\times\mathcal{X}_{U}\times\Theta$.
Assumption 4 (Model exactly matches the data).
For any $\left(X_{O},N_{O},\theta\right)$, there exists some $\left(X_{U},N_{U}\right)$ such that (1) holds. So for
$\mathcal{S}_{\left(f,X_{O},N_{O},\theta\right)}\equiv\left\{\left(X_{U},N_{U}\right)\in\mathcal{X}_{U}\times\mathcal{N}_{U}:f\left(X_{O},X_{U},N_{O},N_{U};\theta\right)=\boldsymbol{0}\right\},$
we have $\mathcal{S}_{\left(f,X_{O},N_{O},\theta\right)}\neq\emptyset$ for all $\left(X_{O},N_{O},\theta\right)\in\mathcal{X}_{O}\times\mathcal{N}_{O}\times\Theta$.
Assumption 3 states that the equilibrium conditions yield a, potentially implicit, unique solution for the endogenous variables $\left(N_{O},N_{U}\right)$ as a function of the exogenous variables $\left(X_{O},X_{U}\right)$. Note that this is a theoretical relationship, and in practice we cannot actually solve these equilibrium conditions for the endogenous variables if there are exogenous unobservables. Assumption 3 will not hold if there are multiple equilibria, meaning that there are at least two equilibria that are not equal up to a multiplicative constant, which I rule out in this paper. Assumption 4 states that the model exactly matches the data, in that for any observables $\left(X_{O},N_{O}\right)$, the unobservables $\left(X_{U},N_{U}\right)$ take on values such that the equilibrium conditions (1) hold exactly. Note that this assumption does not require uniqueness of the unobservables for a given set of observables, and it will often be the case that we cannot separately identify the elements of the unobservables $\left(X_{U},N_{U}\right)$.
A.2 Invertible Models
There is a subset of quantitative spatial models in which there is a one-to-one mapping between the exogenous variables and the endogenous variables. For those models, we are able to invert the model and identify all the exogenous variables from only knowing the endogenous observables. Such models are called invertible in the quantitative spatial literature (Redding and Rossi-Hansberg, 2017). In this case, Assumption 4 will hold with uniqueness. In particular, the exogenous unobservables $X_{U}$ can be recovered from the observables $\left(X_{O},N_{O}\right)$, and there are no endogenous unobservables, so $N_{U}=\left\{\right\}$. One could argue that $X_{U}$ also should be an empty set because we can always recover it, but here I will consider the exogenous unobservables to be variables that are not directly observed in the data.
In these models, we can exactly track the mechanism through which counterfactual exogenous variables map to the levels of counterfactual endogenous variables. The procedure to find these levels has two steps. First, solve for the baseline exogenous unobservables $X_{U}$ from the observables $\left(X_{O},N_{O}\right)$ using Assumption 4 that holds with uniqueness. Then, solve for the levels of counterfactual endogenous variables $N_{O}\hat{N}_{O}$ from the levels of the counterfactual exogenous variables $X_{O}\hat{X}_{O}$ and $X_{U}\hat{X}_{U}$ using Assumption 3.
Remark 5.
In Appendix B.2, I consider a version of the two-country Armington model where trade costs are assumed to be observed. In that case, we can exactly track how measurement errors in various observables affect the levels of the exogenous unobservables, and how these errors ultimately propagate to the levels of the counterfactual endogenous variables.
A.3 Exact Hat Algebra Models
Another important subset of quantitative spatial models contains exact hat algebra models (Costinot and Rodríguez-Clare, 2014). For these models, at least one of $X_{U}$ and $N_{U}$ is non-empty and cannot be recovered from the observables $X_{O}$ and $N_{O}$. We therefore can never identify the counterfactual levels and hence need to shift our focus from levels to changes.
To satisfy Assumption 1 for these models, we need that the endogenous change variables must not depend on the levels of the unobservables. Loosely, in the literature, this is often achieved by having homogeneity or a specific ratio structure in the equilibrium conditions. An example of an exact hat algebra model is the Armington model introduced in Section 2.1.1. There, we cannot identify the baseline levels of the exogenous variables.
Appendix B Details for the Armington Model
B.1 Counterfactual Change Variable System of Equations for Armington
Define the counterfactual level of a scalar variable by $x^{{}^{\prime}}$, so that $x^{{}^{\prime}}=x\hat{x}$. Note that then for this specific counterfactual question, we have $Q_{i}^{{}^{\prime}}=Q_{i}$ for all $i=1,...,n$. Hence, we have
$Y_{i}^{{}^{\prime}}$ $=\sum_{j}\lambda_{ij}^{{}^{\prime}}Y_{j}^{{}^{\prime}}$
$\Rightarrow\hat{Y}_{i}Y_{i}$ $=\sum_{j}\hat{\lambda}_{ij}\lambda_{ij}\hat{Y}_{j}Y_{j}$
for all $i=1,...,n$, and
$\hat{\lambda}_{ij}$ $=\frac{\lambda_{ij}^{{}^{\prime}}}{\lambda_{ij}}$
$=\frac{\left(\tau_{ij}^{{}^{\prime}}Y_{i}^{{}^{\prime}}\right)^{-\varepsilon}\left(Q_{i}^{{}^{\prime}}\right)^{\varepsilon}}{\sum_{k}\left(\tau_{kj}^{{}^{\prime}}Y_{k}^{{}^{\prime}}\right)^{-\varepsilon}\left(Q_{k}^{{}^{\prime}}\right)^{\varepsilon}}\frac{\sum_{\ell}\left(\tau_{\ell j}Y_{\ell}\right)^{-\varepsilon}Q_{\ell}^{\varepsilon}}{\left(\tau_{ij}Y_{i}\right)^{-\varepsilon}Q_{i}^{\varepsilon}}$
$=\frac{\left(\tau_{ij}^{{}^{\prime}}Y_{i}^{{}^{\prime}}\right)^{-\varepsilon}\left(Q_{i}\right)^{\varepsilon}}{\left(\tau_{ij}Y_{i}\right)^{-\varepsilon}Q_{i}^{\varepsilon}}\frac{\sum_{\ell}\left(\tau_{\ell j}Y_{\ell}\right)^{-\varepsilon}Q_{\ell}^{\varepsilon}}{\sum_{k}\left(\tau_{kj}^{{}^{\prime}}Y_{k}^{{}^{\prime}}\right)^{-\varepsilon}\left(Q_{k}^{{}^{\prime}}\right)^{\varepsilon}}$
$=\left(\hat{\tau}_{ij}\hat{Y}_{i}\right)^{-\varepsilon}\frac{\sum_{\ell}\left(\tau_{\ell j}Y_{\ell}\right)^{-\varepsilon}Q_{\ell}^{\varepsilon}}{\sum_{k}\left(\tau_{kj}^{{}^{\prime}}Y_{k}^{{}^{\prime}}\right)^{-\varepsilon}\left(Q_{k}^{{}^{\prime}}\right)^{\varepsilon}}$
$=\frac{\left(\hat{\tau}_{ij}\hat{Y}_{i}\right)^{-\varepsilon}}{\sum_{k}\frac{\left(\tau_{kj}^{{}^{\prime}}Y_{k}^{{}^{\prime}}\right)^{-\varepsilon}\left(Q_{k}^{{}^{\prime}}\right)^{\varepsilon}}{\sum_{\ell}\left(\tau_{\ell j}Y_{\ell}\right)^{-\varepsilon}Q_{\ell}^{\varepsilon}}}$
$=\frac{\left(\hat{\tau}_{ij}\hat{Y}_{i}\right)^{-\varepsilon}}{\sum_{k}\frac{\left(\tau_{kj}^{{}^{\prime}}Y_{k}^{{}^{\prime}}\right)^{-\varepsilon}\left(Q_{k}^{{}^{\prime}}\right)^{\varepsilon}}{\left(\tau_{kj}Y_{k}\right)^{-\varepsilon}\left(Q_{k}\right)^{\varepsilon}}\frac{\left(\tau_{kj}Y_{k}\right)^{-\varepsilon}\left(Q_{k}\right)^{\varepsilon}}{\sum_{\ell}\left(\tau_{\ell j}Y_{\ell}\right)^{-\varepsilon}Q_{\ell}^{\varepsilon}}}$
$=\frac{\left(\hat{\tau}_{ij}\hat{Y}_{i}\right)^{-\varepsilon}}{\sum_{k}\left(\hat{\tau}_{kj}\hat{Y}_{k}\right)^{-\varepsilon}\lambda_{kj}},$
for all $i,j=1,..,n$.
B.2 Running Example: Exactly Identified Version of Armington
For this section again denote the counterfactual level of a scalar variable by $x^{{}^{\prime}}$, so that $x^{{}^{\prime}}=x\hat{x}$. Consider the Armington model where we observe the trade costs:
$X_{O}$ =$=\left\{\tau_{ij}\right\}$
$X_{U}$ =$=\left\{Q_{i}\right\}$
$N_{O}$ =($=\left(\left\{Y_{i}\right\},\left\{\lambda_{ij}\right\}\right)$
[…]
Then, we can identify the endowments or productivities $Q_{i}$ up to a constant:
$\frac{Q_{i}}{Q_{j}}=\frac{Y_{i}}{Y_{j}}\tau_{ij}\left(\frac{\lambda_{ij}}{\lambda_{jj}}\right)^{1/\varepsilon}.$
If there would be no measurement error, then the counterfactual ratio of $Q_{i}$ and $Q_{j}$ resulting from a change in trade costs would be:
$\frac{Q_{i}^{{}^{\prime}}}{Q_{j}^{{}^{\prime}}}=\frac{Y_{i}^{{}^{\prime}}}{Y_{j}^{{}^{\prime}}}\tau_{ij}^{{}^{\prime}}\left(\frac{\lambda_{ij}^{{}^{\prime}}}{\lambda_{jj}^{{}^{\prime}}}\right)^{1/\varepsilon}=\frac{\hat{Y}_{i}Y_{i}}{\hat{Y}_{j}Y_{j}}\hat{\tau}_{ij}\tau_{ij}\left(\frac{\hat{\lambda}_{ij}\lambda_{ij}}{\hat{\lambda}_{jj}\lambda_{jj}}\right)^{1/\varepsilon}=\frac{\hat{Y}_{i}Y_{i}}{\hat{Y}_{j}Y_{j}}\hat{\tau}_{ij}\tau_{ij}\left(\frac{\hat{\lambda}_{ij}F_{ij}}{\hat{\lambda}_{jj}F_{jj}}\right)^{1/\varepsilon}.$
Now, suppose we have homoskedastic independent log-normal errors in bilateral trade costs, income and bilateral trade flows, with measurement error variances $\varsigma_{\tau}^{2}$, $\varsigma_{Y}^{2}$ and $\varsigma_{F}^{2}$ respectively. Then the noisy counterfactual value of the ratio of $Q_{i}$ and $Q_{j}$ will be:
$\frac{\tilde{Q}_{i}^{{}^{\prime}}}{\tilde{Q}_{j}^{{}^{\prime}}}$ $=\frac{\hat{Y}_{i}\tilde{Y}_{i}}{\hat{Y}_{j}\tilde{Y}_{j}}\hat{\tau}_{ij}\tilde{\tau}_{ij}\left(\frac{\hat{\lambda}_{ij}\tilde{F}_{ij}}{\hat{\lambda}_{jj}F_{jj}}\right)^{1/\varepsilon}$
$=\frac{\hat{Y}_{i}Y_{i}e^{\mathcal{N}\left(0,\varsigma_{Y}^{2}\right)}}{\hat{Y}_{j}Y_{j}e^{\mathcal{N}\left(0,\varsigma_{Y}^{2}\right)}}\hat{\tau}_{ij}\tau_{ij}e^{\mathcal{N}\left(0,\varsigma_{\tau}^{2}\right)}\left(\frac{\hat{\lambda}_{ij}F_{ij}e^{\mathcal{N}\left(0,\varsigma_{F}^{2}\right)}}{\hat{\lambda}_{jj}F_{jj}}\right)^{1/\varepsilon}$
$=\frac{Q_{i}^{{}^{\prime}}}{Q_{j}^{{}^{\prime}}}e^{\mathcal{N}\left(0,\varsigma_{\tau}^{2}+2\varsigma_{Y}^{2}+\varsigma_{F}^{2}/\varepsilon^{2}\right)}.$
Here we see that the trade elasticity determines the extend with which the error in trade flows gets propagated to this ratio of $\tilde{Q}_{i}^{{}^{\prime}}$ and $\tilde{Q}_{j}^{{}^{\prime}}$. For $\varepsilon\in\left(0,1\right)$ the error gets inflated and for $\varepsilon>1$ the error gets deflated. The measurement errors in income and in trade flows are not scaled by the trade elasticity. Then, when we want to calculate the counterfactual incomes and expenditure shares for a given change in trade costs, they will also be calculated with noise, and we solve the system of equations in (3) and (4) for $\left\{\tilde{Y}_{i}^{{}^{\prime}}\right\}$ and $\left\{\tilde{\lambda}_{ij}^{{}^{\prime}}\right\}$ using $\left\{\tilde{Q}_{i}^{{}^{\prime}}\right\}$ and $\left\{\tilde{\tau}_{ij}^{{}^{\prime}}\right\}=\left\{\hat{\tau}_{ij}\tilde{\tau}_{ij}\right\}$. This exercise illustrates how the noise in the observables affects the counterfactual endogenous variables through the exogenous unobservables.
Appendix C Alternative method: Adversarial Uncertainty Quantification
The empirical Bayes (EB) approach requires a prior on the true data and a distributional assumption on the measurement errors. A contrasting perspective is taken by what I will call the adversarial approach to uncertainty quantification. This approach borrows insights from the literature on “adversarial perturbations” in machine learning, where small changes to images can lead neural networks to misclassify them with high probability (Moosavi-Dezfooli et al., 2017; Balda, Behboodi, and Mathar, 2019).
C.1 Adversarial Uncertainty Quantification
I transfer these ideas to the setting in this paper and define the perturbed counterfactual function
$\tilde{T}_{\boldsymbol{\eta}}$ $=g_{T}\left(\tilde{D}_{O}+\boldsymbol{\eta}\right),$
where $\tilde{D}_{O}=\left(\tilde{X}_{O},\tilde{N}_{O}\right)$ is the observed data vector and $\boldsymbol{\eta}$ is some perturbation vector. The thought experiment of interest is then: how far are we willing to believe that the observed data is from the true data? That is, denoting with $D_{O}=\left(X_{O},N_{O}\right)$ the true data vector, what are constraints on $\boldsymbol{\eta}$ such that we are almost certain that
$D_{O}\in\left[\tilde{D}_{O}-\boldsymbol{\eta},\tilde{D}_{O}+\boldsymbol{\eta}\right].$
Then, given these constraints on the perturbation $\boldsymbol{\eta}$, we are trying to find the largest deviations in the counterfactual outcome of interest that an adversary can cause. The solutions depend on what kind of norm we use to constrain $\boldsymbol{\eta}$. I opt to use the sup-norm, which corresponds to bounding the adversarial perturbations individually. [^6] To bound the adversarial perturbations jointly, one can use the $L_{2}$-norm. Then, using a linear approximation, the perturbations that an adversary would choose would not depend on the signs of the elements of the Jacobian $\boldsymbol{J}_{T}\left(\tilde{D}_{O}\right)$, but rather on the eigenvector that corresponds to the largest eigenvalue of $\boldsymbol{J}_{T}\left(\tilde{D}_{O}\right)^{{}^{\prime}}\boldsymbol{J}_{T}\left(\tilde{D}_{O}\right)$. The relevant problems to solve then are:
$\boldsymbol{\eta}_{\min}$ $=\arg\max_{\boldsymbol{\eta}:\left|\boldsymbol{\eta}_{j}\right|\leq\boldsymbol{\epsilon}_{j}\forall j}\tilde{T}_{\boldsymbol{0}}-\tilde{T}_{\boldsymbol{\eta}}$
$\boldsymbol{\eta}_{\max}$ $=\arg\max_{\boldsymbol{\eta}:\left|\boldsymbol{\eta}_{j}\right|\leq\boldsymbol{\epsilon}_{j}\forall j}\tilde{T}_{\boldsymbol{\eta}}-\tilde{T}_{\boldsymbol{0}}.$
If the data is low-dimensional, we can brute-force solve these problems. However, as the dimension of the data increases, this becomes increasingly hard. We can use a linear approximation to obtain the following proposition:
Proposition 2.
If we assume that $T$ is approximately linear in the data, and hence $\tilde{T}_{\boldsymbol{\eta}}$ is approximately linear in $\tilde{D}_{O}$ on $\left[\tilde{D}_{O}-\boldsymbol{\epsilon},\tilde{D}_{O}+\boldsymbol{\epsilon}\right]$, we have
$\boldsymbol{\eta}_{\min}$ $\approx-\boldsymbol{\epsilon}\odot\mathrm{sign}\boldsymbol{J}_{T}\left(\tilde{D}_{O}\right)^{{}^{\prime}}$
$\boldsymbol{\eta}_{\max}$ $\approx\boldsymbol{\epsilon}\odot\mathrm{sign}\boldsymbol{J}_{T}\left(\tilde{D}_{O}\right)^{{}^{\prime}},$
where $\odot$ denotes element-wise multiplication, the function $\mathrm{sign}\left(\cdot\right)$ returns the signs of each element of a vector, and $\boldsymbol{J}_{T}\left(\tilde{D}_{O}\right)=\frac{\partial g_{T}}{\partial D_{O}}|_{\tilde{D}_{O}}$ denotes the Jacobian of $g_{T}$ at $\tilde{D}_{O}$.
The proof can be found in Appendix G. This first-order adversarial approach is crude, but it is inexpensive to compute and still provides a useful lower bound. Furthermore, we can get a sense of how close it is to the global bound by comparing the predicted values $\tilde{T}\pm\boldsymbol{J}_{T}\left(\tilde{D}_{O}\right)\left\{\boldsymbol{\epsilon}\odot\mathrm{sign}\boldsymbol{J}_{T}\left(\tilde{D}_{O}\right)^{{}^{\prime}}\right\}$ and the actual values $\tilde{T}_{\pm\boldsymbol{\epsilon}\odot\mathrm{sign}\boldsymbol{J}_{T}\left(\tilde{D}_{O}\right)^{{}^{\prime}}}$ at the first order bounds.
Remark 6.
When we want to simultaneously account for measurement error and estimation error, we have
$\mathbb{P}\left[\hat{k}^{E}\in\left[g_{L_{\hat{k}^{E}}^{\beta}}\left(\tilde{D}_{O}-\boldsymbol{\epsilon}\odot\mathrm{sign}\frac{\partial g_{L_{\hat{k}^{E}}^{\beta}}}{\partial D_{O}}|_{\tilde{D}_{O}}^{{}^{\prime}}\right),g_{U_{\hat{k}^{E}}^{\beta}}\left(\tilde{D}_{O}+\boldsymbol{\epsilon}\odot\mathrm{sign}\frac{\partial g_{U_{\hat{k}^{E}}^{\beta}}}{\partial D_{O}}|_{\tilde{D}_{O}}^{{}^{\prime}}\right)\right]\right]\geq 1-\beta.$
C.2 How to Choose the Bounds on $\boldsymbol{\eta}$
An important question for the adversarial approach to uncertainty quantification is how to choose $\boldsymbol{\epsilon}$. In the applications, I will use the $10\%$ confidence interval of the calibrated measurement error. So if for example we consider normal measurement error with mean 00 and variance $\varsigma^{2}$, the $10\%$ confidence interval will be $\left[-0.13\varsigma,0.13\varsigma\right]$. Researchers could also be agnostic about how large of a confidence interval for the measurement error should be used to choose $\boldsymbol{\epsilon}$. They could then plot the resulting adversarial bounds against the sizes of these confidence intervals and draw a vertical line where they think a valid choice is. I will illustrate such a sensitivity analysis plot for both my applications in Appendices D and E.
Appendix D Details for Application Adao, Costinot, and Donaldson (2017)
D.1 Model Details
In the empirical application of Adao, Costinot, and Donaldson (2017), the authors investigate the effects of China joining the WTO, the so-called China shock. The relevant variables are
$X_{O}$ =$=\left\{Q_{i,t}\right\}$
$X_{U}$ =$=\left\{\tau_{ij,t}\right\}$
$N_{O}$ =($=\left(\left\{\lambda_{ij,t}\right\},\left\{Y_{i,t}\right\}\right)$
$N_{U}$ =$=\left\{P_{i,t}\right\}.$
Here, $Q_{i,t}$ denotes the factor endowment of country $i$ in period $t$, $\tau_{ij,t}$ denotes the trade cost between country $i$ and $j$ in period $t$, $\lambda_{ij,t}$ denotes the expenditure share from country $i$ in country $j$ in period $t$, $Y_{i,t}$ denotes the income of country $i$ in period $t$, and $P_{i,t}$ denotes the factor price of country $i$ in period $t$. Furthermore, $\rho_{i,t}$ denotes the difference between aggregated gross expenditure and gross production in country $i$ in period $t$, which is assumed to stay constant for different counterfactuals. Lastly, $\varepsilon$ denotes the trade elasticity and $\chi_{i}\left(\cdot\right)$ denotes the factor demand system of country $i$.
In Adao, Costinot, and Donaldson (2017), two demand systems are considered, normal CES and “Mixed CES”. I will focus on normal CES, so that
$\lambda_{ij,t}=\chi_{i}\left(\left\{\delta_{ij,t}\right\}\right)=\frac{\exp\left\{\delta_{ij,t}\right\}}{1+\sum_{\ell>1}\exp\left\{\delta_{i\ell,t}\right\}},$
for $\delta_{ij,t}$ some transformation of factor prices. The function $\chi_{i}^{-1}\left(\cdot\right)$ then maps the observed expenditures shares to values of this transformation. The structural parameter $\varepsilon$ is estimated by assuming a model on the unobserved trade costs $\left\{\tau_{ij,t}\right\}$, and is estimated using GMM with as an input the expenditure shares $\left\{\lambda_{ij,t}\right\}$.
The counterfactual question of interest is what the change in China’s welfare is due to joining the WTO. This question is modeled by choosing the counterfactual changes in trade costs, $\hat{\tau}_{ij,t}$, such that Chinese trade costs are brought back to their 1995 levels:
$\hat{\tau}_{ij,t}$ $=\frac{\tau_{ij,95}}{\tau_{ij,t}},\quad\mathrm{if}\>i\>\mathrm{or}\>j\>\mathrm{is}\>\mathrm{China,}$
$\hat{\tau}_{ij,t}$ $=1,\quad\mathrm{otherwise.}$
Welfare is then defined as the percentage change in income that the representative agent in China would be indifferent about accepting instead of the counterfactual change in trade costs from $\left\{\tau_{ij,t}\right\}$ to $\left\{\hat{\tau}_{ij,t}\tau_{ij,t}\right\}$. These changes in China’s welfare $\left\{\hat{W}_{\mathrm{China},t}^{E}\right\}$ can be obtained from first solving for $\left\{\hat{P}_{i,t}\right\}$ using the system of equations
$\sum_{j}\frac{\exp\left\{\chi_{i}^{-1}\left(\left\{\lambda_{ij,t}\right\}\right)-\tilde{\varepsilon}\log\left(\hat{P}_{i,t}\hat{\tau}_{ij,t}\right)\right\}}{1+\sum_{\ell>1}\exp\left\{\chi_{\ell}^{-1}\left(\left\{\lambda_{ij,t}\right\}\right)-\tilde{\varepsilon}\log\left(\hat{P}_{\ell,t}\hat{\tau}_{\ell j,t}\right)\right\}}\left\{\hat{P}_{j,t}\hat{Q}_{j,t}Y_{j,t}+\rho_{j,t}\right\}=\hat{P}_{i,t}\hat{Q}_{i,t}Y_{i,t},$
and then using
$\hat{W}_{i,t}^{E}=\hat{P}_{i,t}\frac{\sum_{\ell}\left[\chi_{\ell}^{-1}\left(\left\{\lambda_{ij,t}\right\}\right)\right]^{-\tilde{\varepsilon}}}{\sum_{\ell}\left[\hat{P}_{\ell,t}\hat{\tau}_{i\ell,t}\chi_{\ell}^{-1}\left(\left\{\lambda_{ij,t}\right\}\right)\right]^{-\tilde{\varepsilon}}}.$
D.2 Estimation details
D.2.1 Calibration Procedure
Combining the prior, measurement error model and the fact that we observe two noisy independent observations per time period for each bilateral trade flow, leads to the following hierarchical model:
$\log\tilde{F}_{ij,t}^{1}$ $=\log F_{ij,t}+\varepsilon_{ij}^{1}$
$\log\tilde{F}_{ij,t}^{2}$ $=\log F_{ij,t}+\varepsilon_{ij}^{2}$
$\log F_{ij,t}$ $=\beta_{t}\log\mathrm{dist}_{ij}+\alpha_{i,t}^{\mathrm{exp}}+\alpha_{j,t}^{\mathrm{imp}}+\eta_{ij},$ (18)
with
$\varepsilon_{ij}^{1},\varepsilon_{ij}^{2}$ $\sim\mathcal{N}\left(0,\varsigma_{ij}^{2}\right)$
$\eta_{ij}$ $\sim\mathcal{N}\left(0,s_{ij}^{2}\right),$
and $\varepsilon_{ij}^{1}$, $\varepsilon_{ij}^{2}$, $\eta_{ij}$ and $\mathrm{dist}_{ij}$ mutually independent. Focusing on variances rather than covariances, we are then interested in estimating the measurement error variances $\left\{\varsigma_{ij}^{2}\right\}$, the prior means $\left\{\log F_{ij,t}\right\}$ and the prior variances $\left\{s_{ij}^{2}\right\}$. To estimate $\varsigma_{ij}^{2}$, note that
$\log\tilde{F}_{ij,t}^{1}-\log\tilde{F}_{ij,t}^{2}\sim\mathcal{N}\left(0,2\varsigma_{ij}^{2}\right)$
and so
$E\left[\left(\log\tilde{F}_{ij,t}^{1}-\log\tilde{F}_{ij,t}^{2}\right)^{2}\right]=2\varsigma_{ij}^{2},$
from which it follows that
$\hat{\varsigma}_{ij}^{2}=\frac{1}{2T}\sum_{t=1}^{T}\left(\log\tilde{F}_{ij,t}^{1}-\log\tilde{F}_{ij,t}^{2}\right)^{2}$
is an unbiased estimator. Towards estimating $\log F_{ij,t}$ and $s_{ij}^{2}$, we can combine the model equations to find:
$\frac{1}{2}\left(\log\tilde{F}_{ij,t}^{1}+\log\tilde{F}_{ij,t}^{2}\right)$ $=\beta_{t}\log\mathrm{dist}_{ij}+\alpha_{i,t}^{\mathrm{exp}}+\alpha_{j,t}^{\mathrm{imp}}+\underbrace{\frac{1}{2}\varepsilon_{ij}^{1}+\frac{1}{2}\varepsilon_{ij}^{2}+\eta_{ij}}_{\equiv\nu_{ij}}.$ (19)
It then follows that an estimator for $\log F_{ij,t}$ is given by the fitted values
$\widehat{\log F_{ij,t}}=\hat{\beta}_{t}\log\mathrm{dist}_{ij}+\hat{\alpha}_{i,t}^{\mathrm{exp}}+\hat{\alpha}_{j,t}^{\mathrm{imp}}.$
Furthermore, since (19) is a valid regression, we can consistently estimate the variance of $\nu_{ij}$. Noting that $\nu_{ij}\sim\mathcal{N}\left(0,\frac{1}{2}\varsigma_{ij}^{2}+s_{ij}^{2}\right)$, it follows that an estimator of $s_{ij}^{2}$ is given by
$\hat{s}_{ij}^{2}$ $=\widehat{\operatorname{Var}}\left(\frac{1}{2}\left(\log\tilde{F}_{ij,t}^{1}+\log\tilde{F}_{ij,t}^{2}\right)-\widehat{\log F_{ij,t}}\right)-\frac{1}{2}\hat{\varsigma}_{ij}^{2}.$
D.2.2 Computational Implementation Details
In preprocessing the mirror trade dataset from Linsi, Burgoon, and Mügge (2023) I made some additional assumptions. Firstly, I only consider data from the period that is considered in Adao, Costinot, and Donaldson (2017). Secondly, I only consider trade flows between countries that the authors of that paper consider. This amounts to aggregating Belgium and Luxembourg, and Estonia and Latvia. All the remaining countries I aggregate to “Rest of World”. Thirdly, when only one of the mirror trade flows is reported, I interpret this as zero measurement error by setting the unknown mirror trade flow equal to the observed one. Relatedly, when both mirror trade flows are not reported, I interpret this as there being no trade, and when one trade flow is zero and the other is substantially larger than zero, I set the zero trade flow equal to the non-zero one. Fourthly, in the mirror trade dataset, the information about trade flows between Taiwan and the three countries Korea, India and Indonesia seems to be missing, and I interpret this as zero trade flows. Lastly, I follow Adao, Costinot, and Donaldson (2017) by setting zero trade flows to 0.0025 (million USD).
When estimating the prior distribution of the true underlying trade flows, I use the distance dataset from Mayer and Zignago (2011). For the distance between countries and the “Rest of World”, I take the average of the distances to all other countries that are considered in Adao, Costinot, and Donaldson (2017).
An important consideration is that there is a substantial difference between the trade flows used in Adao, Costinot, and Donaldson (2017), which come from the World Input Output Dataset (WIOD), and the mirror trade flows from Linsi, Burgoon, and Mügge (2023), which are based on the IMF Direction of Trade Statistics dataset. To overcome this discrepancy, I scale the mirror trade data to make them comparable to the trade flows from WIOD. Since the model Equation (19) has $\frac{1}{2}\left(\log\tilde{F}_{ij,t}^{1}+\log\tilde{F}_{ij,t}^{2}\right)$ as the dependent variable, the scaling I opt to use is $c_{ij,t}\cdot\tilde{F}_{ij,t}^{1}$ and $c_{ij,t}\cdot\tilde{F}_{ij,t}^{2}$ with
$c_{ij,t}=\frac{\tilde{F}_{ij,t}}{\sqrt{\tilde{F}_{ij,t}^{1}\tilde{F}_{ij,t}^{2}}},$
for $\tilde{F}_{ij,t}$ the noisy trade flow as used in Adao, Costinot, and Donaldson (2017). There were also some trade flows in the mirror trade dataset that reported zeros but had a large trade flow in the WIOD. For these trade flows, I set the zero mirror trade data entries equal to the positive WIOD entry. After this replacement exercise, I then use $\exp\left\{\frac{1}{2}\left(\log c_{ij,t}\cdot\tilde{F}_{ij,t}^{1}+\log c_{ij,t}\cdot\tilde{F}_{ij,t}^{2}\right)\right\}$ instead of $\left\{\tilde{F}_{ij,t}\right\}$ as the baseline variable.
Furthermore, for the computational implementation of the bounds that incorporate both estimation error and measurement error, I use the code provided by the authors of Adao, Costinot, and Donaldson (2017) to account for estimation error. That code is somewhat unstable so I make two modifications. Firstly, for some draws of the structural parameter the code was not converging. I opted to ignore these draws when constructing the bounds. Secondly, for the adversarial bounds as in Remark 6 I formally need to take the derivative of the upper and lower bound functions that incorporate estimation error with respect to the observed trade flows, which are $\frac{\partial g_{L_{\hat{k}^{E}}^{\beta}}}{\partial D_{O}}|_{\tilde{D}_{O}}^{{}^{\prime}}$ and $\frac{\partial g_{U_{\hat{k}^{E}}^{\beta}}}{\partial D_{O}}|_{\tilde{D}_{O}}^{{}^{\prime}}$. Since the code that generates these bounds is unstable, I approximate these derivatives by the Jacobian of the point estimate with respect to the observed trade flows, $\frac{\partial g_{\hat{k}}}{\partial D_{O}}|_{\tilde{D}_{O}}^{{}^{\prime}}$.
D.3 Supplementary Analyses
D.3.1 Flat Prior and Homoskedastic Lower Bound for Measurement Error Variances
The first supplementary analysis uses a flat prior and a uniform measurement error variance for all trade flows. This measurement error variance is the average of $\left\{\mathring{\varsigma}_{ij}^{2}\right\}$ from Equation (16), but only considering the countries France, Germany, Great Britain, Japan and the United States, which are all generally considered to be reliable reporters of trade data. This average can hence be interpreted as a lower bound on the measurement error variance, and it equals $\varsigma_{LB}^{2}=0.0048$. The resulting confidence and credible intervals are plotted in Figure 2. Even in this conservative case, the credible set that incorporates both estimation error and measurement error still is large, and for the period 2002-2011 it becomes ambiguous whether joining the WTO had a positive or negative effect on China’s welfare.
D.3.2 Winsorized Measurement Error Variances
The distribution of measurement error variances has a heavy right tail, with the noisiest bilateral trade flow the one from Mexico to Russia with a measurement error variance of 1.14. One might be worried that this heavy tail drives the sensitivity to mismeasurement. Figure 3 replicates Figure 1 but now winsorizing the measurement error variances at 0.2. There are no notable differences between Figures 3 and 1.
D.3.3 Using Adversarial Uncertainty Quantification
Figure 4 replicates Figure 1 but now using the adversarial approach rather than the EB approach. The confidence regions are generally comparable, with the adversarial confidence regions that combine estimation error and measurement error being less extreme than the EB version. For the adversarial approach I use the $10\%$ confidence interval of the measurement error distribution to choose $\boldsymbol{\epsilon}$. This amounts to choosing $\epsilon_{ij}=0.13\varsigma_{ij}$ for all $i,j$, where $\epsilon_{ij}$ is the adversarial bound corresponding to the log trade flow between country $i$ and $j$. Figure 5 depicts the sensitivity analysis plot for the confidence intervals of the measurement error for choosing $\boldsymbol{\epsilon}$, for the changes in welfare in 2011. Interestingly, the lower bounds for different years seems to converge, while the upper bounds diverge.
D.3.4 Probability of Change in Welfare in China Being Larger Than 1
We can learn more from the posterior distributions than just credible sets. It might be of interest what the posterior probability is that the welfare in China increased as a result of the China shock. For the cases with only estimation error or measurement error, this is a trivial exercise. For the case with both estimation error and measurement error I use
$\frac{1}{B}\sum_{b=1}^{B}\min\left\{\max\left\{\frac{U_{b,\hat{W}_{\mathrm{China},t}^{E}}^{\beta}-1}{U_{b,\hat{W}_{\mathrm{China},t}^{E}}^{\beta}-L_{b,\hat{W}_{\mathrm{China},t}^{E}}^{\beta}},0\right\},1\right\}\left(1-\beta\right).$
For the period 1996-2011, the results are plotted in Figure 8. We observe that this probability depends on whether we consider only estimation error, only measurement error, or both. For each of those three cases, the probability seems to be the highest in the period 2006-2007.
D.3.5 Testing Normality Assumption and Gravity Model for the Prior
As outlined in Remark 3, we can check the reasonableness of the normality assumption on the prior by comparing the histogram of the normalized residuals with the probability density function of a standardized normal distribution. The result can be found in Figure 6. It follows that the normality assumption seems reasonable.
Concerning the gravity model, for the year 2011 the regression for the prior mean in Equation (13) has an adjusted R-squared of $0.950$, and the coefficient on log distance is $-0.275$ with a t-statistic of $3.380$. Furthermore, Figure 7 follows Allen and Arkolakis (2018) by plotting a linear and nonparametric fit of log trade flows against log distance, after partitioning out the origin and destination fixed effects. Together, the high adjusted R-squared and the good performance of the linear fit, imply that the gravity model is a reasonable choice for this application.
D.3.6 Variance Decomposition
One might wonder what drives uncertainty in this application, and I use a simple first-order variance decomposition to explore this. For example looking at the change in welfare for China in 2011 due to the China shock, its variance is for 47% explained by the ten most important bilateral trade flows (out of 22,032). Table 3 lists these ten bilateral trade flows.
As expected, many of these important bilateral trade flows involve China and its most important trade partners from 2011. It is however remarkable that Australia and the United States show up multiple times, especially their trade flows with seemingly irrelevant countries for China’s welfare like Slovakia and the Baltic states in the years 1999 and 2000. This can be explained by the fact that in the estimation of the structural parameter, which is the elasticity of substitution, only the data from Australia and the United States are used. This shows that the elasticity of substitution is a strong driver of sensitivity, and that in its estimation some peculiar trade flows play an important role.
Appendix E Details for Application Allen and Arkolakis (2022)
E.1 Model Details
In the empirical application of Allen and Arkolakis (2022), the authors investigate what the returns on investment are of all the highway segments of the US Interstate Highway network. The relevant variables are
$X_{O}$ =$=\left\{\bar{L},\bar{Y}\right\}$
$X_{U}$ =($=\left(\left\{Q_{i}\right\},\left\{A_{i}\right\},\left\{\tau_{ij}\right\}\right)$
$N_{O}$ =($=\left(\left\{F_{ij}\right\},\left\{y_{i}\right\},\left\{\ell_{i}\right\}\right)$
$N_{U}$ =$=\left\{\chi\right\}.$
Here, $\bar{L}$ denotes aggregate labor endowment, $\bar{Y}$ denotes total income in the economy, $Q_{i}$ denotes the productivity of location $i$, $A_{i}$ captures the level of amenities in location $i$, $\tau_{ij}$ denotes the travel cost between locations $i$ and $j$, $F_{ij}$ denotes the traffic flow between locations $i$ and $j$, $y_{i}$ denotes total income of location $i$ as a share of the total income in the economy, $\ell_{i}$ denotes the total labor in location $i$ as a share of the aggregate labor endowment, and $\chi$ captures the (inverse of) the welfare of the economy. The parameter vector is $\theta=\left(\alpha,\beta,\gamma,\delta\right)$, where $\alpha$ and $\beta$ control the strength of the productivity and amenity externalities respectively, $\gamma$ is the shape parameter of the Fréchet distributed idiosyncratic productivity shocks, and $\delta$ governs the strength of traffic congestion.
It is shown in the paper that we can uniquely recover $\left(\hat{y}_{i},\hat{\ell}_{i},\hat{\chi}\right)$ given any change in the underlying infrastructure network $\left\{\hat{\tau}_{ij}\right\}$ and baseline economic activity $\left\{y_{i}\bar{Y}\right\}$, using the system of equations
$\hat{y}_{i}^{\frac{1+\gamma\delta+\gamma}{1+\gamma\delta}}\hat{\ell}_{i}^{\frac{-\theta\left(1+\alpha+\gamma\delta\left(\alpha+\beta\right)\right)}{1+\gamma\delta}}$ $=\hat{\chi}\left(\frac{y_{i}\bar{Y}}{y_{i}\bar{Y}+\sum_{k}F_{ik}}\right)\hat{y}_{i}^{\frac{1+\gamma\delta+\gamma}{1+\gamma\delta}}\hat{\ell}_{i}^{\frac{\gamma\left(\beta-1\right)}{1+\gamma\delta}}$
$\qquad+\sum_{j}\left(\frac{F_{ij}}{y_{i}\bar{Y}+\sum_{k}F_{ik}}\right)\hat{\tau}_{ij}^{\frac{-\gamma}{1+\gamma\delta}}\hat{y}_{j}^{\frac{1+\gamma}{1+\gamma\delta}}\hat{\ell}_{j}^{\frac{-\gamma\left(1+\alpha\right)}{1+\gamma\delta}}$
$\hat{y}_{i}^{\frac{-\gamma\left(1-\delta\right)}{1+\gamma\delta}}\hat{\ell}_{i}^{\frac{\gamma\left(1-\beta-\gamma\delta\left(\alpha+\beta\right)\right)}{1+\gamma\delta}}$ $=\hat{\chi}\left(\frac{y_{i}\bar{Y}}{y_{i}\bar{Y}+\sum_{k}F_{ki}}\right)\hat{y}_{i}^{\frac{-\gamma\left(1-\delta\right)}{1+\gamma\delta}}\hat{\ell}_{i}^{\frac{\gamma\left(\alpha+1\right)}{1+\gamma\delta}}$
$\qquad+\sum_{j}\left(\frac{F_{ji}}{y_{i}\bar{Y}+\sum_{k}F_{ki}}\right)\hat{\tau}_{ij}^{\frac{-\gamma}{1+\gamma\delta}}\hat{y}_{j}^{\frac{-\gamma}{1+\gamma\delta}}\hat{\ell}_{j}^{\frac{\gamma\left(1-\beta\right)}{1+\gamma\delta}}.$
Having obtained $\hat{\chi}$, the change in welfare is then calculated using
$\hat{W}^{E}=\frac{\hat{\chi}^{1/\gamma}}{\bar{L}^{\alpha+\beta}}.$
E.2 Estimation details
E.2.1 Calibration Procedure
The calibration procedure is similar to the one for the application from Adao, Costinot, and Donaldson (2017). Combining the prior and the measurement error model now leads to the following hierarchical model:
$\log\tilde{F}_{ij}$ $=\log F_{ij}+\varepsilon_{ij}$
$\log F_{ij}$ $=\beta\log\mathrm{dist}_{ij}+\alpha_{i}^{\mathrm{orig}}+\alpha_{j}^{\mathrm{dest}}+\eta_{ij},$
with
$\varepsilon_{ij}$ $\sim\mathcal{N}\left(0,0.05\right)$
$\eta_{ij}$ $\sim\mathcal{N}\left(0,s^{2}\right)$
and $\varepsilon_{ij}$, $\eta_{ij}$ and $\mathrm{dist}_{ij}$ mutually independent. We are then interested in estimating the prior means $\left\{\log F_{ij}\right\}$ and the prior variance $\left\{s^{2}\right\}$. Analogously as for Adao, Costinot, and Donaldson (2017), we find
$\widehat{\log F_{ij}}$ $=\hat{\beta}\log\mathrm{dist}_{ij}+\hat{\alpha}_{i}^{\mathrm{orig}}+\hat{\alpha}_{j}^{\mathrm{dest}}$
$\hat{s}^{2}$ $=\widehat{\operatorname{Var}}\left(\log\tilde{F}_{ij}-\hat{\beta}\log\mathrm{dist}_{ij}+\hat{\alpha}_{i}^{\mathrm{orig}}+\hat{\alpha}_{j}^{\mathrm{dest}}\right)-0.05.$
E.2.2 Computational Implementation Details
When I run the code from Allen and Arkolakis (2022), the returns of investment for the links systematically differ slightly from the ones in the paper. I scale my estimates so that the unperturbed estimates align with the ones in the paper. For the procedure that combines measurement error and estimation error, I use the delta method to obtain the frequentist confidence interval that accounts for estimation error rather than a bootstrap procedure for computational reasons.
For the adversarial approach, I only consider measurement error of traffic flows that are adjacent. This is mainly for computational reasons, because calculating the numerical derivative of the returns on investment with respect to a traffic flow is quite computationally expensive. In simulations I observe that these adjacent traffic flows capture most of the variation, in that including the non-adjacent traffic flows will only slightly widen the confidence bounds. Also, for the adversarial approach, similar as for Adao, Costinot, and Donaldson (2017), I approximate the Jacobian at the upper and lower bound by the Jacobian of the point estimate with respect to the traffic flows.
E.3 Supplementary Analyses
E.3.1 Using Adversarial Uncertainty Quantification
The adversarial bounds for the levels and differences of the top three links can be found in Tables 4 and 5. The adversarial approach uses the $10\%$ confidence interval of the measurement error distribution to choose $\epsilon$. The resulting confidence regions are small compared to the ones obtained using the EB approach. Since in Allen and Arkolakis (2022) no estimation error is considered, in Figure 9 I construct the sensitivity analysis plot for the confidence intervals under the assumption of no estimation error. It turns out that in this case the ranking does not change even for large values of the adversarial parameter $\epsilon$.
E.3.2 Probability that Rankings are Reversed
We can learn more from the posterior distributions than just credible sets. It might be of interest what the posterior probability is that the ranking of the three links are reversed. When we consider only measurement error, this probability that the ranking between link 1 and link 2 is reversed equals 0.000 and that the ranking between link 2 and link 3 is reversed equals 0.023. When we consider only estimation error these probabilities change to 0.402 and 0.236 respectively. When we consider both measurement error and estimation error simultaneously, which is calculated using the same approach as in SectionD.3.4, these increase further to 0.600 and 0.584, respectively.
E.3.3 Testing Normality Assumption and Gravity Model for the Prior
We can again check the reasonableness of the normality assumption on the prior as per Remark 3. The result can be found in Figure 10, and it follows that the normality assumption is less reasonable then as for application of Adao, Costinot, and Donaldson (2017).
Concerning the gravity model, the regression for the prior mean in Equation (12) has an adjusted R-squared of $0.9995$, and the coefficient on log distance is $1.003$ with a t-statistic of $1138$. It follows that log distance is an important driver of log traffic flows, but not in a negative way as is common in gravity models. Furthermore, Figure 11 follows Allen and Arkolakis (2018) by plotting a linear and nonparametric fit of log traffic flows against log distance, after partitioning out the origin and destination fixed effects. Together, the high adjusted R-squared and the good performance of the linear fit, imply that the gravity model is a reasonable choice for this application.
Appendix F Misspecification of the Measurement Error Model
Let $L\left(x\right)\propto p\left(\tilde{x}|x\right)$ denote the likelihood function of the noisy data $\left(\tilde{X}_{O},\tilde{N}_{O}\right)$ given the true data $\left(X_{O},N_{O}\right)$. For a given $c>1$ define a density-ratio class of distributions to be the set of all conditional distributions for $\left(\tilde{X}_{O},\tilde{N}_{O}\right)$ with pdf $p$ such that
$p\in\mathcal{R}_{c}=\left\{p\in\mathcal{D}:L\left(x\right)\leq p\left(x\right)\leq c\cdot L\left(x\right)\ \forall x\in\mathcal{X}_{O}\times\mathcal{N}_{O}\right\},$
for $\mathcal{D}$ the set of all pdfs. For any function $g\left(\cdot\right)$, we are then interested in the bounds
$\overline{\mathbb{E}}_{\pi}\left[g\left(X_{O},N_{O}\right)|\tilde{X}_{O},\tilde{N}_{O}\right]$ $=\underset{p\in\mathcal{R}_{c}}{\sup}\mathbb{E}_{\pi}\left[g\left(X_{O},N_{O}\right)|\tilde{X}_{O},\tilde{N}_{O}\right]$
$\underline{\mathbb{E}}_{\pi}\left[g\left(X_{O},N_{O}\right)|\tilde{X}_{O},\tilde{N}_{O}\right]$ $=\underset{p\in\mathcal{R}_{c}}{\inf}\mathbb{E}_{\pi}\left[g\left(X_{O},N_{O}\right)|\tilde{X}_{O},\tilde{N}_{O}\right].$
Denoting with $\pi\left(x\right)$ the pdf of the specified prior-distribution, from Proposition 1 in Geweke and Petrella (1998) we then have the following proposition:
Proposition 3.
$\overline{\mathbb{E}}_{\pi}\left[g\left(X_{O},N_{O}\right)|\tilde{X}_{O},\tilde{N}_{O}\right]$ is the unique solution $t$ to
[…]
and $\underline{\mathbb{E}}_{\pi}\left[g\left(X_{O},N_{O}\right)|\tilde{X}_{O},\tilde{N}_{O}\right]=-\overline{\mathbb{E}}_{\pi}\left[g\left(X_{O},N_{O}\right)|\tilde{X}_{O},\tilde{N}_{O}\right]$.
For the normal-normal model, this leads to the following corollary:
Corollary 1.
In the special case that $L\left(x\right)\propto\phi\left(x;\mu_{m},\Sigma_{m}\right)$ and $\pi\left(x\right)\propto\phi\left(x;\mu_{p},\Sigma_{p}\right)$, $\overline{\mathbb{E}}_{\pi}\left[g\left(X_{O},N_{O}\right)|\tilde{X}_{O},\tilde{N}_{O}\right]$ is the unique solution $t$ to
[…]
with
$\mu=\left(\Sigma_{m}^{-1}+\Sigma_{p}^{-1}\right)^{-1}\left(\Sigma_{m}^{-1}\mu_{m}+\Sigma_{p}^{-1}\mu_{p}\right)\quad\mathrm{and}\quad\Sigma=\left(\Sigma_{m}^{-1}+\Sigma_{p}^{-1}\right)^{-1}.$
We can apply these results for the posterior mean by choosing $g\left(X_{O},N_{O}\right)=g_{T}\left(X_{O},N_{O}\right)$, and we could target $q_{\alpha}$, the $\alpha$-th quantile of the posterior $T|\tilde{X}_{O},\tilde{N}_{O}$, by choosing
$g\left(X_{O},N_{O}\right)=\inf\left\{y\in\mathbb{R}:\mathbb{P}\left[g_{T}\left(X_{O},N_{O}\right)\leq y\right]\geq\alpha\right\}.$
Appendix G Proofs
Proof of Proposition 2
Consider the problem
$\boldsymbol{\eta}_{\max}=\arg\max_{\boldsymbol{\eta}:\left|\boldsymbol{\eta}_{j}\right|\leq\boldsymbol{\epsilon}_{j}\forall j}\tilde{T}_{\boldsymbol{\eta}}-\tilde{T}_{\boldsymbol{0}}.$
Note that if $\tilde{T}_{\boldsymbol{\eta}}$ is approximately linear in $\tilde{D}_{O}$ on $\left[\tilde{D}_{O}-\boldsymbol{\epsilon},\tilde{D}_{O}+\boldsymbol{\epsilon}\right]$, we have
$\tilde{T}_{\boldsymbol{\eta}}\approx\tilde{T}_{\boldsymbol{0}}+\boldsymbol{J}_{T}\left(\tilde{D}_{O}\right)\boldsymbol{\eta}.$
Hence, to maximize $\tilde{T}_{\boldsymbol{\eta}}-\tilde{T}_{\boldsymbol{0}}$ we will always choose $\boldsymbol{\eta}_{j}$ at the boundary of $\left[-\boldsymbol{\epsilon}_{j},\boldsymbol{\epsilon}_{j}\right]$, where the choice of which boundary is governed by the sign of the $j$-th element of $\boldsymbol{J}_{T}\left(\tilde{D}_{O}\right)$.
Appendix H Appendix Tables and Figures
Year Exporter Importer Percentage of variance explained
2011 Rest of World China 29.94%
2011 Japan China 3.35%
2011 Korea China 2.44%
1999 Estonia-Latvia Australia 2.21%
1999 Estonia-Latvia United States 2.21%
2011 Germany China 1.60%
2011 Taiwan China 1.44%
2000 Slovakia Australia 1.33%
2000 Slovakia United States 1.33%
2010 Lithuania Australia 1.23%
Table 3: The ten bilateral trade flows that explain the largest share of the variance in the change in welfare for China in 2011 due to the China shock.
Link 1 Link 2 Link 3
Point estimate 10.43 9.54 7.31
Only est. error [5.31, 15.55] [4.69, 14.38] [3.62, 11.00]
Only meas. error, adversarial [10.17,10.76] [9.31,9.74] [6.87,7.89]
Est. and meas. error, adversarial [5.53, 17.10] [5.07,15.48] [3.72, 12.56]
Table 4: Adversarial uncertainty quantification for the three links from Allen and Arkolakis (2022) with the highest return on investment. Link 1 is Kingsport-Bristol (TN-VA) to Johnson City (TN), link 2 is Greensboro-High Point (NC) to Winston-Salem (NC) and link 3 is Rochester (NY) to Batavia (NY).
Link 1-Link 2 Link 2-Link 3
Point estimate 0.89 2.23
Only est. error [-6.16,7.94] [-3.86,8.32]
Only meas. error, adversarial [0.79,1.24] [1.98,2.42]
Est. and meas. error, adversarial [-5.56, 9.09] [-169.87, 9.11]
Table 5: Adversarial uncertainty quantification for the differences between the three links from Allen and Arkolakis (2022) with the highest return on investment. Link 1 is Kingsport-Bristol (TN-VA) to Johnson City (TN), link 2 is Greensboro-High Point (NC) to Winston-Salem (NC) and link 3 is Rochester (NY) to Batavia (NY).
Figure 2: EB uncertainty quantification for homoskedastic normal shocks to $\left\{\log F_{ij,t}\right\}$ for the change in China’s welfare due to the China shock. The solid blue line is the estimate as reported in Adao, Costinot, and Donaldson (2017).
Figure 3: EB uncertainty quantification for winsorized heteroskedastic normal shocks to $\left\{\log F_{ij,t}\right\}$ for the change in China’s welfare due to the China shock. The solid blue line is the estimate as reported in Adao, Costinot, and Donaldson (2017).
Figure 4: Adversarial uncertainty quantification for heteroskedastic normal shocks to $\left\{\log F_{ij,t}\right\}$ for the change in China’s welfare due to the China shock. The solid blue line is the estimate as reported in Adao, Costinot, and Donaldson (2017).
Figure 5: Sensitivity analysis plot for the change in China’s welfare in 2011 due to the China shock. At zero, the bounds collapse to the estimates as reported in Adao, Costinot, and Donaldson (2017). The vertical line indicates the choice used throughout this paper.
Figure 6: Plot to compare the normalized residuals with the probability density function of a standardized normal distribution to check whether the normality assumption for the prior is reasonable for Adao, Costinot, and Donaldson (2017).
Figure 7: Plot that follows Allen and Arkolakis (2018) to check whether the gravity model is reasonable for log trade flows in 2011 from Adao, Costinot, and Donaldson (2017).
Figure 8: Probability that the change in welfare in China is larger than 1, following Adao, Costinot, and Donaldson (2017).
Figure 9: Sensitivity analysis plot for levels and differences of the three US highway links with the highest return on investment. At zero, the bounds collapse to the estimates as reported in Allen and Arkolakis (2022). The vertical line indicates the choice used throughout this paper.
Figure 10: Plot to compare the normalized residuals with the probability density function of a standardized normal distribution to check whether the normality assumption for the prior is reasonable for Allen and Arkolakis (2022).
Figure 11: Plot that follows Allen and Arkolakis (2018) to check whether the gravity model is reasonable for log traffic flows from Allen and Arkolakis (2022).